### v1

# On the correlation between Architectural Smells and Static Analysis Warnings

Abstract
Background. Software quality assurance is essential during software development and maintenance. Static Analysis Tools (SAT) are widely used for assessing code quality. Architectural smells are becoming more daunting to address and evaluate among quality issues.
Objective. We aim to understand the relationships between Static Analysis Warnings (SAW) and Architectural Smells (AS) to guide developers/maintainers in focusing their effort on SAWs more prone to co-occurring with AS.
Method. We performed an empirical study on 103 Java projects totaling 72 million LOC belonging to projects from a vast set of domains, and 785 SAW detected by four SAT, Checkstyle, Findbugs, PMD, SonarQube, and 4 architectural smells detected by ARCAN tool. We analyzed how SAWs influence AS presence. Finally, we proposed an AS remediation effort prioritization based on SAW severity and SAW proneness to specific ASs.
Results. Our study reveals a moderate correlation between SAWs and ASs. Different combinations of SATs and SAWs significantly affect AS occurrence, with certain SAWs more likely to co-occur with specific ASs. Conversely, 33.79% of SAWs act as ”healthy carriers,” not associated with any ASs.
Conclusion. Practitioners can ignore about a third of SAWs and focus on those most likely to be associated with ASs. Prioritizing AS remediation based on SAW severity or SAW proneness to specific ASs results in effective rankings like those based on AS severity.
keywords:
Static Analysis Tool , Software Quality Warnings , Architectural Smells , Technical Debt
journal: Journal of Systems and Software
1 Introduction
The risk of service degradation and disruption for critical services is exponentially increasing due to the pervasive nature of computing and the widespread adoption of digital services in all areas of daily life [1]. To create a common definition for quality issues, the literature provides the concept of smells [2], referring to quality issues in code [3]. These smells are usually categorized as code smells, low level, coding issues [4, 2], and, recently, in architectural smells, higher abstract issues that concern the software architecture and the software components [5]. More specifically, architectural smells (AS) are “architectural decisions that negatively impact system internal quality” [6], which can occur simply by “applying a design solution in an inappropriate context, mixing design fragments that have undesirable emergent behaviors, or applying design abstractions at the wrong level of granularity” [6]. High-level quality issues are more difficult to spot given the non-direct, and evident connection, to low-level coding issues [7]. Therefore, recent studies stressed the role of AS in architectural degradation, thus hastening architectural technical debt accumulation  [8, 9, 10].
A popular technique to detect quality issues is to leverage Static Analysis Tools (SAT) [11, 12] which analyze the source code without executing it (thus static analyzers), and produce a report containing all the violations, i.e., Static Analysis Warnings (SAW) that the code present. More specifically, SAWs, refer to code violations with a lower-level granularity, compared to AS. Therefore, SAWs do not consider the software architecture of the system under development, commonly focusing on class or method levels [13]. While multiple studies successfully exploited SATs to remediate common quality issues [13, 14, 3, 15, 16], only a few studies focused on developing SATs to detect ASs [5]. Moreover, some studies have investigated the interrelations between code anomalies [17, 18] or code smells [19] and AS [17, 18]. However, to our knowledge, no previous study has analyzed the correlations between SAW and AS. Furthermore, to ensure the safety and stability of services, during development and maintenance, AS remediation should be prioritized [9, 10]. Hence, the portion of code affected by AS is the primary remediation target to prevent critical quality issues.
Therefore, the need to investigate how practitioners and researchers can exploit SAWs to detect and remediate AS is evident. Hence we designed and performed a large empirical study on the possible correlations between four AS and SAW detected by four tools largely adopted by developers [13].
Hence, this study assesses the correlation between SAW and higher-level AS. We did not consider correlations between defects and SAW, as these have been extensively studied in the literature. Additionally, we investigate the correlation of SAW and AS, and how SAW can be utilized to prioritize AS. Discovering a correlation between AS and SAW would enable the use of existing SAT as an effective tool for detecting issues in software architecture. For practitioners, linking SAW to architectural smells can reduce the need for refactoring, thereby saving both cost and effort [20, 21]. Therefore, we can summarize our main contributions as follows:
  • We performed the first, according to our knowledge, large-scale study, i.e., 112 projects, on the correlation of ASs and SAWs. Other studies have considered correlations with more general architectural problems [18, 17] or the correlation between AS and code smells [19].
  • We investigated the contribution of the most correlated SAWs with ASs.
  • We proposed and evaluated AS remediation effort prioritization via SAWs severity.
Our study focused on 112 Java projects from the Qualitas Corpus dataset (QCD) [22], totaling 72 million lines of code analysis and 66,036 packages, 808,208 classes, and 88,460 interfaces examined across various contexts, including development platforms, multimedia, gaming, middleware, compilers, and testing frameworks. We considered a large set of warnings (1142) and four AS based on dependency issues that can critically influence the quality of a software project and its progressive architecture degradation [23].
Our findings reveal a moderate correlation between SAWs and ASs. Moreover, different combinations of SATs and SAWs significantly influence the occurrence of ASs. Therefore specific SAWs are more prone to co-occur with specific ASs. Conversely, we also identified SAWs acting as “healthy carriers,” i.e., not co-occurring with any ASs. Healthy carriers represent 33.79% of our dataset. Our study shows that practitioners can safely drop a third of the SAW in our context and focus on the most AS-prone to address AS without prior AS severity or discovery knowledge effectively. Our study highlighted that prioritizing AS remediation efforts based on SAW severity, or SAW proneness to specific AS with a limited inspection window, results in rankings similar to the optimal one, i.e., ordered by AS severity.
Therefore, we suggest that SAWs can influence AS’s presence. Prioritizing SAWs based on occurrence probability effectively guides remediation. Future research should improve the detection of AS-prone SAWs and explore deeper connections between SAWs and AS for better understanding and prioritization of remediation efforts.
Paper Structure. Section 2 describes the background to which our paper is based. In Section 3, we present the study design. In Section 4, we show the results obtained and discuss them in Section 5. Section 6 focuses on threats to the validity of our study. Section 7 discusses related work and in Section 8, we draw conclusions and outline future research directions.
2 Background
2.1 Static Analysis Tool
This section introduces the tools we employed in our study and the motivation for their selection. SAT analyzes software without running it by inspecting the source code, thus discovering potential quality issues in the code base [8]. Recently, the ease-to-useness of SATs and the subsequent inclusion in CI practices [24] are boosting their popularity. Among the SAT available, we selected the most widely used according to state of the art in detecting SAW [11, 13] such as Sonarqube, Checkstyle, Findbugs, and PMD. While according to architectural issues, our study focuses on architectural smells (AS). Therefore, we also included ARCAN based on our previous experience [5]. In Table 1, we summarized the characteristics of each of the following SATs.
SAW Checkstyle Findbugs PMD SonarQube
Warnings Types 14 types 9 types 8 types 3 types
Total Warnings 173 424 305 413
Types Annotations, Block Checks, Class Design, Coding, Headers, Imports, Javac Comments, Metrics, Miscellaneous, Modifiers, Naming Conventions, Regexp, Size Violations, Whitespace Bad practice, Correctness, Experimental, Internationalization, Malicious code vulnerability, Multithreaded correctness, Performance, Security, Dodgy code Best Practices, Code Style, Design, Documentation, Error Prone, Multithreading, Performance, Security Bugs, Code Smells, Vulnerabilities
Severity Levels Error, Ignore, Info, Warning 1-20 ranking: Scariest, Scary, Troubling, Concern Change required, Change suggested, Change optional, Change highly recommended, Change highly optional Blocker, Critical, Major, Minor, Info
Table 1: Summary of Static Analysis Tools Warning (SAW)
More specifically, regarding SATs, Checkstyle [^1] https://checkstyle.org is designed to asses Java code quality. The tool uses “rules” according to a set of checks for analyzing the code base. Checkstyle has two predefined rule configurations: Google Java Style and Sun Java Style. Static violations of the checks are grouped under two severity levels: error and warning.
FindBugs [^2] http://findbugs.sourceforge.net examine Java bytecode. The tool detects bug patterns which are caused by, but not limited to, difficult language features and misunderstood API features. Bug patterns are ranked on an ordinal scale from one to 20. Rank 1-4 is the scariest group, rank 5-9 is the scary group, rank 10-14 is the troubling group, and rank 15-20 is the concern group.
PMD [^3] https://pmd.github.io/ can inspect eight programming languages, including Java. PMD uses a set of rules to assess code quality. There are a total of 33 different rule-set configurations for Java projects. Rules are classified into 8 categories. The tool measures rule violations on an ordinal scale of one to five. One is associated with the most severe violations, and 5 is the least severe.
SonarQube [^4] http://www.sonarsource.org/ is one of the most common SATs for detecting code quality issues. SonarQube measures several aspects of the source code, such as several lines of code and code complexity, and verifies code compliance with a specific set of coding rules defined for most common programming languages. Violations of coding rules are reported as issues in SonarQube. The estimated time to remove these issues, i.e., remediation effort, is used to compute the remediation cost. SonarQube analysis encompasses quality aspects such as reliability, maintainability, and security. Reliability rules, also named Bugs, create quality issues that “represent something wrong in the code” and that will soon be reflected in a bug. Code smells are considered “maintainability related issues” in the code that decrease code readability and modifiability. It is important to note that the term “code smells” adopted in SonarQube does not refer to the commonly known code smells term defined by Fowler et al. [4], but to a different set of rules.
On the other hand, we employed ARCAN to detect different AS based on dependency issues by computing different metrics proposed by Martin [25], such as those related to instability issues. We refer to software instability as the inability to make changes without impacting the entire system or a large part.
The tool relies on graph database technology to perform graph queries, which lets on higher scalability during the detection process and management of many different kinds of dependencies. The detailed detection techniques for AS have been described in previous works [5]. Moreover, the tool allows the computation of an architectural debt [26] index based on the number of AS detected in a project and the criticality of the AS, evaluated according to each type of AS. The tool’s results were validated on ten open source systems and two industry projects with a high precision value of 100% in the results and 63% recall [5]. Moreover, the results of ARCAN were validated using practitioners’ feedback on four industry projects [27]. In our study, we have considered the following AS detected by the tool:
  1. 1.
    Unstable Dependency (UD): describes a subsystem (component) that depends on other less stable subsystems than itself [28]. This may cause a ripple effect of changes in the system [5]. Detected in packages.
  2. 2.
    Hub-Like Dependency (HD): This smell arises when an abstraction has (outgoing and ingoing) dependencies on a large number of other abstractions [29]. Detected in classes and packages.
  3. 3.
    Cyclic Dependency (CD): refers to a subsystem (component) that is involved in a chain of relations that break the desirable cyclic nature of a subsystem’s dependency structure. The subsystems involved in a dependency cycle can hardly be released, maintained, or reused in isolation. Detected in classes and packages. The Cyclic Dependency AS is detected according to different shapes [30] as described in [5].
We focused on the above AS based on dependency issues since dependencies are very important in software architecture. Components with many dependencies can be considered more critical and expensive to maintain. Moreover, we defined as “healthy” the instances without AS, i.e., a software entity affected by one or more SAWs but without any of the ASs considered in our study.
2.2 Effort Aware Metrics
Çarka et al. [31] investigated effort-aware metrics (EAM) to improve defect prioritization. In our study, we take inspiration from the definition of PofB and PopT to create three different prioritization rankers for AS. We then evaluate our ranker in a way that is similar to the interpretation of the PofB results. Therefore, we briefly discuss EAMs; according to Çarka et al. [31], two different EAMs exist: normalized by size or not by size. The most known not-normalized EAM is called PofB [32, 33, 34, 35] PofB, or Proportion of Defects Identified by the top xof the codebase, is a measure derived from a prediction model’s rankings. It signifies the percentage of defective entities discovered when analyzing the initial xof the code. Higher PofB indicates more effective rankings, thereby enhancing testing support. For instance, a PofB10 of 30% implies that 30% of defective entities were detected by analyzing 10% of the codebase according to the method’s ranking.
Comparing a prediction model’s ranking with a perfect one sheds light on its performance. Mende et al. [36], drawing inspiration from Arisholm et al. [37], introduced ”Popt” to gauge this accuracy. It assesses how the prediction model deviates from perfection while outperforming random rankings. Popt quantifies the area ${\Delta opt}$ between the ideal and predicted models. In the optimal scenario, instances are ordered by decreasing fault density, while in the predicted model, they’re ordered by decreasing predicted defectiveness. The equation for Popt, ${1-\Delta opt}$, indicates a larger value denotes a smaller gap between the ideal and predicted models [38].
Popt and PofB are distinct metrics, each capturing different facets of model accuracy. They employ different ranking methods: Popt ranks entities based on bug density (bug probability divided by entity size), while PofB ranks them according to bug probability. Consequently, classifiers ranked by Popt and PofB may differ. Popt, grounded in bug density, offers a more realistic perspective than PofB, which relies solely on probability. However, interpreting Popt is more challenging; a classifier with double the Popt value doesn’t necessarily yield double the benefits to the user.
3 Empirical Study Design
This section presents the empirical study’s goal, research questions, metrics, and hypotheses. Moreover, we describe the study context, the data collection, and the data analysis procedures. We designed and conducted our empirical study according to the guidelines defined by Wohlin et al. [39].
3.1 Goal, Research Questions, Metrics, and Hypothesis
We formalized the goal of this study according to the Goal Question Metric (GQM) approach [40] as follows:
The goal of our empirical study is to investigate the relationship between architectural smells and static analysis warnings in the context of software quality assurance. Our perspective is of practitioners and researchers seeking to assess the relationship between high-level architectural issues and low-level code quality problems.
Based on our goal, we formulate the following Research Questions (RQs)
RQ1. Are SAWs and AS correlated?
Investigating the correlation between AS and SAW enables us to assess the relationship between high-level architectural issues and low-level code quality problems. We aim to determine whether ASs are correlated with SAWs and the strength of this correlation. To address RQ1, we collected SAW and AS presence per software package. To assess statistical significance, we conjectured one hypothesis ($H_{1}$) as follows:
  • $H_{11}$: There is a statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells.
Hence, we defined the null hypothesis ($H_{0}$) as follows:
  • $H_{01}$: There is no statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells.
We then tested whether the SAW influenced the presence of ASs computing Spearman’s $\rho$ correlation coefficient [41] between the total number of specific SAWs and ASs, averaged per software package (see Section 3.4).
Moreover, different static analyzer warnings can lead to specific architectural smells. Therefore, assessing whether a specific SAW’s presence is a possible indicator of specific AS presence is essential. Therefore, we ask:
RQ2. Do SAWs influence AS presence?
In the long run, architectural smell can silently deteriorate complex software architecture. Hence, we investigate the characteristics of SAW affecting software projects and leading to AS to identify common traits among projects prone to AS and assess further factors contributing to AS emergence and persistence. To address RQ2 we analyzed the co-occurrence of SAW and ASs. Let $\sigma$ be a SAW and $\alpha$ be an AS, and $\eta$ be a software entity. We state that $\sigma$ co-occur with $\alpha$, if given $\eta$ such that $\sigma$ and $\alpha$ affect contemporary $\eta$. To assess statistically significance, we conjectured two hypotheses ($H_{2}$ and $H_{3}$ ) as follows:
  • $H_{12}$: There is a statistically significant difference in the co-occurrence of pairs of SAW and AS.
  • $H_{13}$: There is a statistically significant difference in the co-occurrence of SAT-specific SAWs’ and AS presence.
Hence, we defined the null hypotheses ($H_{0}$) as follows:
  • $H_{02}$: There is no statistically significant difference in the co-occurrence of pairs of SAW and AS.
  • $H_{03}$: There is no statistically significant difference in the co-occurrence of SAT-specific SAWs’ and AS presence.
We computed the co-occurrence of each SAW and AS. For each project package, we counted the number of healthy CDs, UDs, and HLs (and combinations). We tested $H_{2}$ and $H_{3}$ with Wilcoxon Paired Signed Ranked Test (WPT) (see Section 3.4).
Finally, assessing the correlation between SAW and AS and understanding how this relationship influences the presence of AS in software projects merely scratches the surface. Indeed, it falls short, considering practitioners’ limited time for remediation efforts. Therefore, prioritizing these efforts becomes paramount. Thus, we ask:
RQ3. Are SAWs effective for AS remediation effort prioritization?
Effort prioritization is essential to avoid wasting resources on noncritical tasks [42]. No prior study has investigated AS prioritization via SAW severity. According to the “no free lunch theorem”, currently, there is no default severity for code quality checks universally accepted by our community [43]. Therefore, in our study, we investigate exploiting SAW characteristics, such as SonarQube rules severity and cost estimation, the prioritization effort for AS severity according to Arcelli Fontana et al. [5]. We referred to recent empirical studies to assess the accepted severity of AS [44, 45, 46]. More specifically, Sas and Avgeriou [46] propose, on average, an ordinal scale of one to ten, where one is the least severe AS and ten is the highest. Hence, we adopted the average value of CD (5), UD (7), and HL (9). Moreover, regarding SAW severity, we compiled a comprehensive dataset in our previous study [2] focused on SonarQube. Without the risk of generalizability issue [43], we focused this RQ on SonarQube-specific SAWs. To address RQ3 we quantify the probability of a SAW inducing an AS ($\mathcal{P}$) as the relative frequency at which the particular SAW contributed to the occurrence of an AS. To assess statistical significance, we conjectured one hypothesis ($H_{4}$) as follows:
  • $H_{14}$: There is a statistically significant difference in the prioritized effort between the three rankers.
Hence, we defined the null hypothesis ($H_{0}$) as follows:
  • $H_{04}$: There is no statistically significant difference in the prioritized effort between the three rankers
We computed $\mathcal{P}$ for each SAT, SAW, and AS. We ranked the SAWs based on their severity [2], $\mathcal{P}$, and AS severity [46]. We tested $H_{4}$ with WPT (see Section 3.4).
3.2 Study Context
We selected projects from the Qualitas Corpus (QCD) collection of software systems [47], using the compiled version of the QCD [22]. Figure 1 shows the word cloud composed by the names of the 112 java projects in the dataset crafted by Tempero et al. [22]. Table 2 details metrics on the QCD.
Figure 1: Qualitas Corpus WordCloud [22].
Metric Value
Number of Projects 112
Lines of Code (LOC) $>$ 18 million
Packages 16,509
Classes Analyzed 202,052
Interfaces Analyzed 22,115
Methods Analyzed 464,893
Contexts Covered IDEs, SDKs, databases, 3D/graphics/media, diagram/visualization libraries and tools, games, middlewares, parsers/generators/make tools, programming language compilers, testing libraries, and tools.
Available Metrics LOC, NOP, NOCL, NOI, NOM, NOA, NORM, PAR, NSM, NSA. CK Metrics: WMC, DIT, NOC, LCOM HS. - MLOC, SIX, VG, NBD, RMD. - CA, CE, I, A.
Table 2: Summary of Qulitas Corpus Metrics
3.3 Study setup and data collection
Figure 2: Data Collection Workflow
This section presents our data collection methodology. Figure 2 presents the study workflow. Our data collection comprised three steps:
  1. 1.
    We analyzed the QCD with four SATs and averaged the results for project packages;
  2. 2.
    We separately analyzed QCD with the Arcan tool;
  3. 3.
    To get our study results, we analyzed the finalized dataset, comprising all SAW detected by the three SATs and the AS detected by the Arcan tool grouped by software package.
We analyzed the entire data set (112 Java projects) from the Qualitas Corpus [47]. However, we obtained results for only 103 projects because the computational time for the remaining ones would have been excessively long.
3.4 Data Analysis
We considered the classes affected by at least an AS and one static analysis warning at the class or package levels. One or more static analysis warnings could infect the same set of classes for each AS. This is because an AS may involve more than one class, while static analysis warnings involve only one class. In the case of projects not infected by static analysis warnings or AS, we did not consider them for the analysis.
To answer RQ1, we collected the QCD projects and ran the SAT and ARCAN. We merged the outputs of the tool. In our context, the dependent variable is the AS, and the independent variables are the SAW. To compute the correlation coefficients, we need to assess the distributions of our variable. Therefore we conjectured one hypothesis ($H_{\mathcal{N}}$) as follows:
  • $H_{1\mathcal{N}}$: The data do not follow a normal distribution.
Hence, we defined the null hypothesis ($H_{0}$) as follows:
  • $H_{0\mathcal{N}}$: The data follow a normal distribution.
We tested $H_{\mathcal{N}}$ with the Anderson-Darling (AD) test [48]. The AD test assesses whether data samples derive from a specific probability distribution, such as the normal distribution. AD measures the difference between the sample data and the expected values from the tested distribution. More specifically, it evaluates differences in the cumulative distribution function (CDF) between the observed data and the hypothesized distribution [48].
The AD test and the Shapiro-Wilk (SW) test [49] are both statistical tests used to assess the normality of data.
The SW test focuses on the correlation between the observed data and the expected values under a normal distribution, emphasizing the smallest and largest values in the dataset. Therefore, according to Mishra et al. [50], SW is more appropriate when targeting small datasets ($\leq 50$ samples). On the other hand, the AD test considers a wider range of values, including those in the middle of the distribution, providing a more sensitive evaluation of normality, especially for larger sample sizes than our own. Therefore, we preferred AD over SW [51]. AD is also considered one of the most powerful statistical tools for detecting most deviations from normality [51, 52].
Test results allowed us to reject the null hypothesis for each test case, $H_{0\mathcal{N}}$, asserting that the data was not normally distributed. Therefore, we chose Spearman’s $\rho$ [41] to compute the correlation in place of Pearsons, which requires a normal distribution [53]. The non-parametric Spearman’s rank correlation coefficient, denoted by $\rho$, measures statistical dependence between two variables without necessarily assuming a linear relationship or even normality of those variables. Spearman’s Rho measures the strength and direction of the monotonic relationship between the ranks of paired data points. Table 3 presents $\rho$ values interpretation according to Dancey and Reidy [54].
Interpretation Perfect Strong Moderate Weak Zero
Correlation + 1 + 0.9 +0.8 + 0.7 +0.6 + 0.5 + 0.4 + 0.3 + 0.2 + 0.1 0
- 1 - 0.9 - 0.8 - 0.7 - 0.6 - 0.5 - 0.4 - 0.3 - 0.2 - 0.1 0
Table 3: Spearman’s $\rho$ interpretation (RQ1)
To answer RQ2, we computed Spearman’s $\rho$ for each SAW-AS combination. We used the interquartile ratio (IQR) method [55] to select the top correlated SAW for each SAT and AS, i.e., we computed the quartiles. We selected only the SAW with their Spearman’s $\rho$ value higher or equal to the third quartile (Q3), thus selecting the top 25%. We computed the $\mathcal{P}$ as follows: Let $\Sigma_{\text{SAT}_{i}}$ represent the set of Static Analysis Warnings specific to a particular Static Analysis Tool $\text{SAT}_{i}$. Then, $\sigma\in\Sigma_{\text{SAT}_{i}}$ signifies that $\sigma$ belongs to the set of SAWs specific to $\text{SAT}_{i}$; thus $\sigma_{j}$ is a specific SAW and $\alpha\in AS$ represents a specific AS.. Therefore, we define the relative frequency $\mathcal{P}(\sigma_{j},\alpha)$ as:
$\mathcal{P}(\sigma_{j},\alpha)=\frac{|\sigma_{j}\text{ in }\alpha|}{|\sigma\text{ in }\alpha|}$
Where:
  • The numerator represents the number of times the specific SAW, $\sigma_{j}\in\Sigma_{\text{SAT}_{i}}$, appears within the given AS, $\alpha$.
  • The denominator represents the total count of all SAWs belonging to a specific SAT, $\sigma$, in the given AS, $\alpha$.
More specifically, we computed $\mathcal{P}$ per each possible combination of AS. Therefore our data computed a $\mathcal{P}$-value for the 3 ASs, the healthy instances, and for combinations of ASs. We tested $H_{2}$, and $H_{3}$ via the WPT [56], which is a non-parametric statistical test that compares two related samples or paired data. WPT uses the absolute difference between the two observations to classify and then compare the sum of the positive and negative differences. The test statistic is the lowest of both. We selected WPT to test $H_{2}$ and $H_{3}$ because the data was not normally distributed; hence, we used it instead of the paired t-test, which assumes a normal data distribution.
Moreover, due to the high number of statistical tests performed, we needed to apply p-value correction. Correcting p-values is essential in multiple hypothesis testing to control the increased risk of Type I errors. When multiple tests are conducted simultaneously, the probability of incorrectly rejecting at least one true null hypothesis increases. A classical p-value correction technique is the Bonferroni correction. However, Bonferroni is known to offer a conservative correction [57]. Therefore, we decided to use the Benjamini & Hochberg (BH) p-value correction [58], which uses the false discovery rate (FDR) when performing multiple hypothesis tests as follows [59]:
  1. 1.
    Rank the p-values: Sort the p-values obtained from multiple tests in ascending order, denoted as $p_{(1)}\leq p_{(2)}\ldots\leq p_{(m)}$ for their respectively defined hypotheses $H_{(1)},H_{(2)},\ldots,H_{(m)}$.
  2. 2.
    Assign ranks: Assign each p-value a rank, $i$, with $p_{(1)}$ being the smallest p-value and $p_{(m)}$ being the largest.
  3. 3.
    Calculate the corrected p-values: For each p-value $p_{(i)}$, calculate the BH-corrected p-value using the formula:
    $q_{(i)}=\min_{j\in\{1,\ldots,m\}}\left\{\frac{m\cdot p_{(j)}}{j},1\right\}$
    where $m$ is the total number of tests, $q_{(i)}$ is the corrected probability for the hypothesis at moment $i$ and $p_{(j)}$ given the order rank $j\in\{1,\ldots,m\}$ of p-values, which stands as the probability at the rank position $j$.
  4. 4.
    Determine correction index: The detailed procedure identified the largest index $k$ for which BH holds and rejects null hypotheses as follows:
    $k=\max\left\{i\in{1,\ldots,m}\mid p_{(i)}\leq i\frac{\alpha}{m}\right\}$
    which is automatically computed with existing analytical software.
  5. 5.
    Corrected hypothesis testing: Compare the corrected p-values $q_{(i)}$ from each hypothesis $H_{(1)},H_{(2)},\ldots,H_{(m)}$ against the $k$ index, that is:
    $p_{(i)}\leq i\frac{\alpha}{m}$
    to the desired critical value, $\alpha$ (e.g., 0.05), to determine which hypotheses to reject.
Since we performed the BH correction on the p-values of the defined hypotheses, we maintained the standard value of the critical value $\alpha$ on 0.05, hence gaining power on controlling for error Type I and Type II [59].
To answer RQ3, we leverage our definition of $\mathcal{P}$, the default severity associated with the specific SAW, and the severity associated with the specific AS to build three different rankers. The ranking idea stemmed from previous research by Çarka et al. [31], in which PofB and PopT metrics were normalized and used to evaluate classifier prediction concerning inspection effort (see Section 2.2).
Following the definition of PofB and Popt, we designed three rankers. More specifically, we ordered the data according to:
  • SAW severity: we used SonarQube severity associated with each SAW.
  • $\mathcal{P}$: we used $\mathcal{P}$ and created separate rankings one for each AS combination.
  • AS severity: we used the known AS severity to order the data.
We note that the AS severity-based order is a special case; by design, that ranker is the optimal one because it uses the AS severity. RQ3 aims to use the other two rankers to prioritize those SAWs that are more prone to induce ASs. Therefore, the optima ranker is needed to compare the results of the other two. According to the interpretation of PofB and PopT, we selected the first xof the data from the three rankings. Therefore, for each selected percentage of data, we computed how many Critical, High, and Medium AS-prone SAWs we would capture.
3.5 Replicability
We provide a replication package containing the raw data, including the instructions for the SAT execution, the list of projects infected by SAW and AS, and the full statistical test results for both the normality hypothesis and for the RQ’s hypotheses  [^5] https://zenodo.org/doi/10.5281/zenodo.11366846.
4 Results
In this section, we reported the results we obtained to answer our RQs.
4.1 Static Analysis Warnings and Architectural Smell Correlation (RQ1)
We tested $H_{1}$ with Spearman’s $\rho$. We excluded all cross-correlations among SAWs and focused on the correlations between SAW and AS. Figure 3 presents the distribution of Spearman’s $\rho$ averaged for SAW by AS and SAT. We can reject the null hypothesis for each of the 663 cases. Hence, we can affirm that there is a statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells. More specifically, SAWs were weak to moderately correlated to AS for each AS across SAT.
Figure 3: Spearman’s $\rho$ average distribution for SAW by AS and SAT (RQ1)
4.2 Static Analysis Warnings inducing Architectural Smell (RQ2)
We tested $H_{2}$ with WPT. We rejected the null hypothesis for 243,258 out of 275,380 cases, i.e., 88%. Therefore, we can, on average, affirm that there is a statistically significant difference in the co-occurrence of pairs of SAW and AS. More specifically, Figure 4 presents the distribution of $H_{2}$ test results. It is worth noticing that in the case of UD, there was almost an equal probability of rejecting or accepting the null hypothesis across SAWs. Moreover, we note that there is an equal probability of rejecting the null hypothesis in the case of Healthy cases as well as those affected by all of the three ASs.
Furthermore, we tested $H_{3}$ with WPT. We can reject the null hypothesis in 30 out of 42 cases, i.e., 72% of the cases. Table 4 shows the test cases for which we could not reject the null hypothesis. Therefore, we can affirm that there is a statistically significant difference in the co-occurrence of SAT-specific SAWs and ASs presence.
Figure 4: Distribution of $H_{2}$ test results (RQ2)
Table 4: Wilcoxon All Pair Test, i.e., H03 p-value (RQ2)
AS Healthy CD CD CD CD CD
SAT PMD PMD FindBugs PMD SonarQube SonarQube
Compared with CheckStyle CheckStyle CheckStyle FindBugs CheckStyle FindBugs
p-value 0,76 0,25 0,44 0,85 0,58 0,20
AS CD UD UD + CD HL UD + HL + CD UD + HL + CD
SAT SonarQube FindBugs PMD PMD FindBugs PMD
Compared with PMD CheckStyle CheckStyle CheckStyle CheckStyle CheckStyle
p-value 0,02 1,00 0,14 0,40 0,03 0,28
More specifically, Table 5 presents the top five SAW co-occurring with specific AS. We selected the top five SAWs that co-occur the most for each AS, including the “healthy carriers”. We note that the most co-occurring SAWs belong to FindBugs. Similarly, FindBugs also has the most “healthy carriers” co-occurring SAWs. Moreover, where three ASs affect a single package, the FindBugs SAW co-occurs each time. Futhermore, the least co-occurring AS is UD.
Table 5: The top five SAW co-occurring with specific AS
SAT SAW AS % co-occurring
FindBugs AT_OPERATION_SEQUENCE_ON_CONCURRENT_ABSTRACTION UD + HL + CD 100
FindBugs DMI_EMPTY_DB_PASSWORD 100
FindBugs DMI_ENTRY_SETS_MAY_REUSE_ENTRY_OBJECTS 100
FindBugs DMI_RANDOM_USED_ONLY_ONCE 100
FindBugs EQ_COMPARING_CLASS_NAMES 100
PMD UseStringBufferForStringAppends UD + HL 6,132
PMD AssignmentInOperand 4,856
PMD code_smells:spaghetti_code 3,448
PMD UseCollectionIsEmpty 2,435
PMD BooleanGetMethodName 1,904
SonarQube squid:ClassCyclomaticComplexity UD + CD 3,418
PMD AvoidPrefixingMethodParameters 3,222
SonarQube squid:S1197 2,300
FindBugs BC_UNCONFIRMED_CAST_OF_RETURN_VALUE 2,270
FindBugs BC_UNCONFIRMED_CAST 2,178
SonarQube S2386 UD 0,429
SonarQube S1444 0,294
SonarQube S1141 0,267
SonarQube S1118 0,242
SonarQube S1226 0,223
FindBugs DMI_COLLECTION_OF_URLS HL 40,000
FindBugs OBL_UNSATISFIED_OBLIGATION 33,871
CheckStyle NoWhitespaceAfterCheck 31,748
FindBugs EQ_DOESNT_OVERRIDE_EQUALS 21,739
PMD LongVariable 18,457
FindBugs NS_NON_SHORT_CIRCUIT CD 88,235
FindBugs ICAST_INTEGER_MULTIPLY_CAST_TO_LONG 87,500
FindBugs RV_CHECK_FOR_POSITIVE_INDEXOF 80,000
FindBugs DM_GC 78,571
FindBugs VA_FORMAT_STRING_USES_NEWLINE 75,000
FindBugs NM_METHOD_NAMING_CONVENTION Healthy 89,865
FindBugs RI_REDUNDANT_INTERFACES 86,029
FindBugs SIO_SUPERFLUOUS_INSTANCEOF 70,130
FindBugs EQ_CHECK_FOR_OPERAND_NOT_COMPATIBLE_WITH_THIS 68,421
PMD CompareObjectsWithEquals 66,212
4.3 Architectural Smell Prioritization (RQ3)
Figure 5 presents the distribution of $\mathcal{P}$ averaged among SAW across SAT per AS. Therefore, on average, Figure 5 presents the percentage of occurrences of SAWs related to the occurrence of one or more ASs or their absence (i.e., healthy). On average, SAWs belonging to SonarQube are more prone to raise warnings on healthy packages and classes. Most SAWs co-occur with CD AS and have a similar presence among the SATs. Moreover, second to CD, SAWs co-occurred with all three ASs. In this context, FindBugs’ SAWs exhibit an average higher co-occurrence.
Figure 5: Distribution of $\mathcal{P}$ averaged among SAW across SAT per AS (RQ2)
Figure 6 presents the SAW prioritization based on the three rankings, i.e., SAW Severity-based, $\mathcal{P}$-based, and optimal. We computed how many medium, high, or critical AS could be discovered by inspecting the first xof data ranked by one of the three methods. According to its design, the optimal ranking prioritized the SAW co-occurring with high and critical AS. The SAW severity-based and UD+CD rankers demonstrate the most balanced performance, while the ranking based on SAW co-occurring with CD favors moderately severe AS.
Finally, we tested $H_{4}$ with the WPT. We could reject the null hypothesis in all cases for high and critical AS severities, in which the optimal ranker was compared with the SAW-severity-based and the $\mathcal{P}$ based rankers. Conversely, we could reject the hypothesis when comparing the SAW-severity-based and the $\mathcal{P}$ based rankers but only the specific case when comparing the UD + HL + CD proneness based on $\mathcal{P}$. Futhermore, when the difference was statically significant, the gains were in favor of the optimal rankers, as expected.
We included the full table and the graphical representation of the gains among the approaches in the replication package for space constraint.
Moreover, we can not reject the null hypothesis for the medium risk for each ranker. Therefore, the optimal ranker behaves similarly to the other two rankers.
Hence, we can affirm that there is a statistically significant difference in the prioritized effort between the three rankers.
Nevertheless, based on the results, we can affirm that SAW prioritization based on $\mathcal{P}$ can effectively guide AS remediation effort.
Figure 6: Comparison of Inspection Effort Analysis based on SAW Severity, $\mathcal{P}$, and AS Severity.
5 Discussion
This section discusses the results and presents lessons learned based on them.
Our findings reveal a moderate correlation between SAWs and ASs. However, different combinations of SATs and SAWs significantly influence the occurrence of ASs. Thus, our study’s key takeaway is that SAWs have a statistically discernible influence on ASs, despite the moderate correlation.
Furthermore, specific ASs show varying degrees of susceptibility to SAWs. For instance, as illustrated in Figure 4, UD’s acceptance or rejection of $H_{03}$ was almost equally probable. Likewise, SAWs that consistently fail to trigger ASs, i.e., Healthy, consistently reject the null hypothesis. As a result, this SAWs, acting as “healthy carriers”, can be safely discarded when prioritizing AS remediation efforts.
More specifically, the prevalence of FindBugs SAWs among the top co-occurring SAWs indicates that FindBugs might be more effective at finding issues related to ASs. This shows that FindBugs is effective at finding certain AS patterns. Thus, future research efforts should focus on the difference between FindBugs SAW and the other SATs SAWs. Thus grasping the characteristics that make FindBugs SAWs co-occur with all three ASs and with the “healthy carriers”.
Futhermore, future research efforts should focus on identifying UD AS, being the least co-occurring AS may suggest that these issues are not sufficiently characterized by the considered SAWs and SATs.
Therefore, by avoiding wasting remediation time on healthy carries, regarding $RQ_{3}$, our study highlighted that prioritizing AS remediation efforts based on SAW severity, or SAW proneness to specific AS with a limited inspection window, results in similar rankings to the optimal one. Healthy carriers represent 33.79% of our dataset. Hence we can safely drop a third of the SAW, focusing on the most AS-prone to effectively address AS without prior AS-severity or discovery knowledge.
6 Threats to Validity
In this section, we discuss the threats to the validity of our study. We categorized the threats into Construct, Internal, External, and Conclusion, following the guidelines defined by Wohlin et al. [39].
Construct Validity. Construct validity concerns how our measurements reflect what we claim to measure [39].
Our specific design choices may impact our results, including our measurement process and data filtering. To address this threat, we based our choice on past studies [5, 22, 60, 46] and used well-established guidelines in designing our methodology [61, 40]. More specifically, we used four static analysis tools considered to be the most adopted by developers to detect software quality warnings [13] and ARCAN to detect the four architectural smells.
Nevertheless, we acknowledge that the analyzed SAW does not represent all the possible SAWs currently available in the state-of-the-art. We mitigated this threat by selecting the most popular SAT. Nonetheless, future works will focus on extending the current selection of SAW. Similarly, we analyzed only three AS due to the lack of AS detection tools. Future works will focus on expanding the AS detectable by ARCAN to mitigate this threat.
Internal Validity. Internal Validity is the extent to which an experimental design accurately identifies a cause-and-effect relationship between variables [39]. Our study relies on a large-scale analysis of 103 popular Java projects in the Qualitas Corpus dataset. However, they can potentially be biased from their project selection. We address this threat by designing our inclusion criteria based on past studies [60, 47, 22].
External Validity. External validity concerns how the research elements (subjects, artifacts, etc.) represent actual elements [39]. External validity concerns how the research elements (subjects, artifacts, etc.) represent actual elements [39]. The analyzed project comes from the GitHub public repository. Mining versioning systems, particularly GitHub, also threaten external validity. More specifically, GitHub’s user base predominantly comprises developers and contributors to open-source projects, potentially skewing findings towards this specific demographic. In our context, this is not a real threat because it is the study’s focus. Moreover, the dynamic nature of GitHub, with frequent updates, forks, and merges, poses challenges in ensuring the stability and consistency of data over time. We addressed this issue by providing the raw data in our replication package. Furthermore, the accessibility of GitHub data is subject to various permissions and restrictions set by project owners, potentially hindering reproducibility and transparency in research. We addressed this research, limiting our data collection to public repositories. Nonetheless, the vast diversity of projects analyzed in our study presents a significant diversity in terms of programming languages, project sizes, and development methodologies, thus aiding the generalizability of results.
Conclusion Validity focuses on how we draw conclusions based on the design of the case study, methodology, and observed results [39]. The distribution of our data determines the type of statistical tools we can use to test our hypothesis. We tested the data to asses if it departed from normality with the AD test, and we could reject the null hypothesis for each test. Therefore we selected the WT test instead of the t-test as Spearman’s $\rho$ in place of Pearsons. Furthermore, statistical tests threaten the conclusion’s validity regarding the appropriateness of statistical tests and procedures, such as assumption violation, multiple comparisons, and Type I or Type II errors. We address this issue using WT instead of the t-test due to the rejection of the normal data distributions. Moreover, we applied the HB correction [62] to balance type 1 and type 2 errors well. We also set our alpha to 0.01, hence reducing it from the standard value of 0.05 due to the many statistical tests we performed.
Spearman’s $\rho$ is a non-parametric measure of the strength and direction of a correlation between two variables, but due to multiple factors, this approach may hinder the conclusion’s validity. For instance, the correlation coefficient is unreliable with a small sample size and cannot accurately represent the relationship between variables. Moreover, Spearman assumes a linear monotonic correlation between the variables; if the assumption were to prove invalid or not applicable, Spearman’s correlation might not accurately measure the strength and direction of the association. We have mitigated this issue by ensuring a representative sample and checking the data distribution. We could mitigate this issue by using other non-parametric measures of correlation. Still, our analysis found that Spearman’s $\rho$ was more suitable than other correlation measures, such as Kendall’s $\tau$, because it can effectively handle tied observations in the data. On the other hand, Kendall’s $\tau$ relies on concordant and discordant pairs of observations.
7 Related Work
This section discusses works related to our study regarding SATs and AS.
7.1 Static Analysis Tools
In this section, we report the relevant work on SAT focusing on their usage, warnings, and the detected issues [63, 64, 65]. SATs increasing in popularity [11, 66] given their ease-to-use [24]. Recently researchers focus on analyzing SAT usage  [67, 68, 69, 70] and effectiveness [13]. For instance, SATs effectively improve bug prediction models [71]. Developers use SAT to inspect software codebases, identify bugs [72, 73], evaluate code quality [74, 75, 76], and address coding issues [77, 78, 15].
Recent studies has explored SAT’s application across various programming languages [65], their configurations [65], and their evolution within software systems [79, 80, 77]. However, fewer studies have focused on the impact of SAT on software quality [79, 81, 15, 82, 83].
Moreover, the research community also focused on prioritization strategies for addressing warnings. For instance, researchers proposed prioritization based on removal times [84], or developer engagement in remediating specific warnings [76]. Moreover, studies have compared estimated warning resolution times with actual developer efforts [80, 85, 86, 87].
Researchers have extensively studied SonarQube regarding the impact of SAW on software quality. They investigated the fault-proneness of SonarQube rules [79, 15, 88] and their susceptibility to changes [81, 15]. Findings indicate that SonarQube rules increase change-proneness at the class level, while correlations between faults and rules can reduce fault-proneness [79, 15].
Similarly, researchers have examined PMD warnings concerning their impact on pull request acceptance. Results show quality warnings do not affect pull request acceptance [82]. A comprehensive study [83] on warning trends in OSS projects using PMD highlighted that large-scale changes in SAT warnings often result from coding style modifications, such as braces and naming conventions. The presence of PMD in build configurations positively impacts defect density, suggesting an improvement in external quality.
Recently, Wright et al. [89] investigated why developers use SAT, revealing new tool requirements such as systems for recommending warnings based on developer expertise and collaborative interfaces for warning analysis. Conversely, Johnson et al. [74] examined why developers do not widely use SAT tools, identifying false positives and the presentation of warnings as significant barriers. Furthermore, Yang et al. [90] explored enhancing the utility of static code warning tools through data mining algorithms that filter out commonly ignored warnings, using FindBugs as a case study.
7.2 Architectural Smells
AS have been less thoroughly investigated than code smells. While numerous studies have explored correlations between code smells, research into the relationship between code smells and AS has been limited. One study found a minimal correlation between AS and code smells, suggesting that AS cannot be reliably inferred from code smells [19].
The connection between code anomalies, similar to code smells, and AS has also been studied with mixed results. Some findings indicated that many detected code anomalies did not correspond to architectural issues [17], while other results showed that over 80% of architectural problems were linked to code smells [18]. Specific code smells like Long Method, and God Class were consistently associated with architectural problems.
Oizumi et al. [91] investigate the aggregation of code anomalies and their relationship to AS and show that a single anomaly can signal an architectural problem. Specifically, 50% of syntactic aggregations and 80% of semantic aggregations are linked to design issues [92, 91]. A study does not find a strong correlation between AS and architectural degradation, outlining that AS cannot explain architectural degradation alone [93], while Sas et al. [94] highlights the effects of AS on long-term system maintainability and evolvability, supported by interview excerpts. Practitioners highlight which parts of the AS analysis offer actionable insights for planning refactoring activities.
Moreover, Mumtaz et al. [95] shows that community smells, particularly the Missing Links smell, are related to design smells. They discuss specific refactoring techniques that concurrently address community and design smells, managing social and technical issues together.
Finally, Fontana et al. [19] on the correlation between technical debt indices based on code-level issues from SonarQube revealed no correlation, suggesting differing impacts based on the features considered [26]. Although several tools have been developed for detecting architectural debt, including DV8, Designite, Jarchitect, and others, which focus on AS in various programming languages [96, 97, 98], no prior studies have examined correlations between AS and SAW as we have addressed in this paper.
8 Conclusion
Our results show a statistically significant positive correlation between SAWs and AS, showing weak to moderate correlations across different tools and SAWs and AS in 661 out of 663 tests. In terms of co-occurrence, we found a statistically significant difference in the pairs of SAWs and ASs, with 88% of the Wilcoxon All-Pair Ranking Test cases rejecting the null hypothesis. Specifically, we showed that SAWs can influence the presence of ASs. When prioritizing architectural smells, our findings indicate that using a prioritization method based on the probability of SAW occurrences ($\mathcal{P}$) can effectively guide remediation efforts, revealing a statistically significant difference between the three ranking methods tested. Our proposed approach to SAW prioritization highlights that, in the absence of AS indication, SAW severity and our empirically computed $\mathcal{P}$ together with the possibility of dropping a third of the data to inspect, proved to be effective ASs remediation effort prioritization.
Future research efforts should enhance SAT capabilities to detect the most AS-prone SAWs more accurately, improving AS remediation prioritization. Additionally, research should aim to deepen the analysis of SAWs’ AS-proneness to better trace AS back to SAWs. This will help understand the motivations or root causes linking specific SAWs to specific AS.

### v2

@@
 Abstract
-Background.
+Context.
 Software quality assurance is essential during software development and maintenance.
@@
 Architectural smells are becoming more daunting to address and evaluate among quality issues.
-Objective.
-We aim to understand the relationships between Static Analysis Warnings (SAW) and Architectural Smells (AS) to guide developers/maintainers in focusing their effort on SAWs more prone to co-occurring with AS.
-Method.
-We performed an empirical study on 103 Java projects totaling 72 million LOC belonging to projects from a vast set of domains, and 785 SAW detected by four SAT, Checkstyle, Findbugs, PMD, SonarQube, and 4 architectural smells detected by ARCAN tool.
-We analyzed how SAWs influence AS presence.
-Finally, we proposed an AS remediation effort prioritization based on SAW severity and SAW proneness to specific ASs.
+Purpose.
+We aim to understand the relationships between Static Analysis Warnings (“warnings”) and Architectural Smells (“smell”) to guide developers/maintainers in focusing their effort on warnings more prone to co-occurring with smell.
+Methods.
+We performed an empirical study on 103 Java projects totaling 72 million LOC belonging to projects from a vast set of domains, and 785 warnings were detected by three SAT, Checkstyle, Findbugs, PMD, SonarQube, and 4 architectural smells were detected by the ARCAN tool.
+We analyzed how warnings influence smell presence.
+Finally, we proposed an smell remediation effort prioritization based on warning severity and warning proneness to specific smells.
 Results.
-Our study reveals a moderate correlation between SAWs and ASs.
-Different combinations of SATs and SAWs significantly affect AS occurrence, with certain SAWs more likely to co-occur with specific ASs.
-Conversely, 33.79% of SAWs act as ”healthy carriers,” not associated with any ASs.
+Our study reveals a moderate correlation between warnings and smells.
+Different combinations of SATs and warnings significantly affect smell occurrence, with certain warnings more likely to co-occur with specific smells.
+Conversely, 33.79% of warnings are “non-co-occurring” with any of the smells in our dataset.
+This provides an early indicator for potential architectural concerns before resource-intensive architectural analysis is performed.
 Conclusion.
-Practitioners can ignore about a third of SAWs and focus on those most likely to be associated with ASs.
-Prioritizing AS remediation based on SAW severity or SAW proneness to specific ASs results in effective rankings like those based on AS severity.
+Practitioners can ignore about a third of warnings and focus on those most likely to be associated with smells.
+Prioritizing smell remediation based on warning severity or warning proneness to specific smells results in effective rankings like those based on smell severity.
+While not a substitute for specialized tools like ARCAN, warning-based prioritization provides a pragmatic bridge between low-level warnings and high-level architectural issues, particularly useful in contexts lacking full architectural visibility.
 keywords:
 Static Analysis Tool, Software Quality Warnings, Architectural Smells, Technical Debt
-journal: Journal of Systems and Software
 1 Introduction
-The risk of service degradation and disruption for critical services is exponentially increasing due to the pervasive nature of computing and the widespread adoption of digital services in all areas of daily life [1].
+The risk of architectural erosion [98, 71] and disruption for critical services is exponentially increasing due to the pervasive nature of computing and the widespread adoption of digital services in all areas of daily life [21].
 To create a common definition for quality issues, the literature provides the concept of smells [44], referring to quality issues in code [89].
-These smells are usually categorized as code smells, low level, coding issues [4, 2], and, recently, in architectural smells, higher abstract issues that concern the software architecture and the software components [5].
-More specifically, architectural smells (AS) are “architectural decisions that negatively impact system internal quality” [6], which can occur simply by “applying a design solution in an inappropriate context, mixing design fragments that have undesirable emergent behaviors, or applying design abstractions at the wrong level of granularity” [6].
-High-level quality issues are more difficult to spot given the non-direct, and evident connection, to low-level coding issues [7].
-Therefore, recent studies stressed the role of AS in architectural degradation, thus hastening architectural technical debt accumulation  [8, 9, 10].
-A popular technique to detect quality issues is to leverage Static Analysis Tools (SAT) [11, 12] which analyze the source code without executing it (thus static analyzers), and produce a report containing all the violations, i.e., Static Analysis Warnings (SAW) that the code present.
-More specifically, SAWs, refer to code violations with a lower-level granularity, compared to AS.
-Therefore, SAWs do not consider the software architecture of the system under development, commonly focusing on class or method levels [13].
-While multiple studies successfully exploited SATs to remediate common quality issues [13, 14, 3, 15, 16], only a few studies focused on developing SATs to detect ASs [5].
-Moreover, some studies have investigated the interrelations between code anomalies [17, 18] or code smells [19] and AS [17, 18].
-However, to our knowledge, no previous study has analyzed the correlations between SAW and AS.
-Furthermore, to ensure the safety and stability of services, during development and maintenance, AS remediation should be prioritized [9, 10].
-Hence, the portion of code affected by AS is the primary remediation target to prevent critical quality issues.
-Therefore, the need to investigate how practitioners and researchers can exploit SAWs to detect and remediate AS is evident.
-Hence we designed and performed a large empirical study on the possible correlations between four AS and SAW detected by four tools largely adopted by developers [13].
-Hence, this study assesses the correlation between SAW and higher-level AS.
-We did not consider correlations between defects and SAW, as these have been extensively studied in the literature.
-Additionally, we investigate the correlation of SAW and AS, and how SAW can be utilized to prioritize AS.
-Discovering a correlation between AS and SAW would enable the use of existing SAT as an effective tool for detecting issues in software architecture.
-For practitioners, linking SAW to architectural smells can reduce the need for refactoring, thereby saving both cost and effort [20, 21].
+These smells are usually categorized as code smells, low-level coding issues [30, 44], and, recently, architectural smells, higher-level abstract issues that concern the software architecture and the software components [3].
+A popular technique to detect quality issues is to leverage Static Analysis Tools (SAT) [96, 5] which analyze the source code without executing it (thus static analyzers), and produce a report containing all the violations, i.e., Static Analysis Warnings (from now on “warnings”) that the code presents.
+More specifically, warnings refer to code violations with a lower-level granularity compared to smells.
+Therefore, warnings do not consider the software architecture of the system under development, commonly focusing on class or method levels [24].
+Despite the large body of work on smells detection techniques using architectural reconstruction, graph-based models, or clustering heuristics [60], these methods are often resource-intensive, demand explicit architectural views, or require specialized tooling that may not be readily integrated into developers’ daily workflows.
+In contrast, warnings are pervasive in software engineering practice.
+They are automatically produced, tightly coupled with development environments, and consulted regularly by practitioners.
+This makes warnings a pragmatic and low-barrier data source, yet their potential to act smells early indicators of architectural erosion remains largely underexplored.
+While multiple studies successfully exploited SATs to remediate common quality issues [24, 69, 89, 48, 2], only a few studies focused on developing SATs to detect smells [3].
+Moreover, some studies have investigated the interrelations between code anomalies [54, 53] or code smells [28] and smells [54, 53].
+However, to our knowledge, no previous study has analyzed the correlations between warning and smells.
+Furthermore, to ensure the safety and stability of services, during development and maintenance, smell remediation should be prioritized [66, 74].
+Hence, the portion of code affected by smells is the primary remediation target to prevent critical quality issues.
+This work investigates whether warnings, widely available and frequently ignored, can serve as useful signals to identify or prioritize architectural remediation efforts.
+By doing so, we aim to bridge the gap between commonly available code-level signals and hard-to-detect architectural issues, providing practical insights for both researchers and practitioners.
+Therefore, the need to investigate how practitioners and researchers can exploit warnings to detect and remediate smells is evident.
+Hence, we designed and performed a large empirical study on the possible correlations between four smells and warning detected by three tools largely adopted by developers [24].
+Our study assesses the correlation between warning and higher-level smells.
+We did not consider correlations between defects and warnings, smells, these have been extensively studied in the literature.
+Specifically, we investigate the correlation between warning and smells and how warning can prioritize smells.
+Past studies have struggled to show an obvious relationship between warnings and smells [72].
+Nonetheless, discovering a correlation between smells and warnings would enable the use of existing SAT smells as an effective tool for detecting issues in software architecture.
+Importantly, this approach requires no additional tooling or modeling effort, making it highly attractive in industrial contexts.
+For practitioners, linking warning to architectural smells can reduce the need for refactoring, thereby saving both cost and effort [17, 16].
 Therefore, we can summarize our main contributions as follows:
-  • We performed the first, according to our knowledge, large-scale study, i.e., 112 projects, on the correlation of ASs and SAWs.
-Other studies have considered correlations with more general architectural problems [18, 17] or the correlation between AS and code smells [19].
-  • We investigated the contribution of the most correlated SAWs with ASs.
-  • We proposed and evaluated AS remediation effort prioritization via SAWs severity.
-Our study focused on 112 Java projects from the Qualitas Corpus dataset (QCD) [22], totaling 72 million lines of code analysis and 66,036 packages, 808,208 classes, and 88,460 interfaces examined across various contexts, including development platforms, multimedia, gaming, middleware, compilers, and testing frameworks.
-We considered a large set of warnings (1142) and four AS based on dependency issues that can critically influence the quality of a software project and its progressive architecture degradation [23].
-Our findings reveal a moderate correlation between SAWs and ASs.
-Moreover, different combinations of SATs and SAWs significantly influence the occurrence of ASs.
-Therefore specific SAWs are more prone to co-occur with specific ASs.
-Conversely, we also identified SAWs acting as “healthy carriers,” i.e., not co-occurring with any ASs.
-Healthy carriers represent 33.79% of our dataset.
-Our study shows that practitioners can safely drop a third of the SAW in our context and focus on the most AS-prone to address AS without prior AS severity or discovery knowledge effectively.
-Our study highlighted that prioritizing AS remediation efforts based on SAW severity, or SAW proneness to specific AS with a limited inspection window, results in rankings similar to the optimal one, i.e., ordered by AS severity.
-Therefore, we suggest that SAWs can influence AS’s presence.
-Prioritizing SAWs based on occurrence probability effectively guides remediation.
-Future research should improve the detection of AS-prone SAWs and explore deeper connections between SAWs and AS for better understanding and prioritization of remediation efforts.
+  • To our knowledge, we performed the first large-scale study, i.e., 103 projects, on the correlation of smells and warnings.
+Other studies have considered correlations with more general architectural problems [53, 54] or the correlation between smells and code smells [28].
+  • We investigated the contribution of the most correlated warnings with smells.
+  • We proposed and evaluated smells remediation effort prioritization via warnings severity.
+Our study focused on 103 Java projects from the Qualitas Corpus dataset (QCD) [90], totaling 72 million lines of code analysis and 66,036 packages, 808,208 classes, and 88,460 interfaces examined across various contexts, including development platforms, multimedia, gaming, middleware, compilers, and testing frameworks.
+We considered a large set of warnings (1142) and four smells based on dependency issues that can critically influence the quality of a software project and its progressive architecture erosion [27, 51, 61].
+Contrary to past studies, our findings reveal a moderate, yet statistically significant, correlation between warnings and smells.
+Moreover, different combinations of SATs and warnings significantly influence the occurrence of smells.
+Therefore, specific warnings are more prone to co-occur with specific smells.
+Conversely, we also identified that warnings non-co-occurring with any smells represent 33.79% of our dataset.
+Our study shows that practitioners can safely drop a third of the warnings in our context and focus on the most smell-prone areas to address smells without prior smell severity or discovery knowledge effectively.
+Our study highlighted that prioritizing smell remediation efforts based on warning severity, or warning proneness to specific smells with a limited inspection window, results in rankings similar to the optimal one, i.e., ordered by smell severity.
+Therefore, we suggest that warnings can moderately influence the presence of smells.
+Prioritizing warnings based on occurrence probability effectively guides remediation.
+Future research should improve the detection of smell-prone warnings and explore deeper connections between warnings and smells for better understanding and prioritization of remediation efforts.
 Paper Structure.
@@
 Section 7 discusses related work and in Section 8, we draw conclusions and outline future research directions.
+Definitions: In the remainder of this paper, we refer to static analysis warnings as ‘warnings’ and architectural smells as ‘smells’.
 2 Background
-2.1 Static Analysis Tool
+This section introduces the static analysis tools used, the types of architectural smells considered, and the effort-aware metrics applied to evaluate prioritization.
+2.1 Static Analysis Tools and Warnings
 This section introduces the tools we employed in our study and the motivation for their selection.
 SAT analyzes software without running it by inspecting the source code, thus discovering potential quality issues in the code base [20].
-Recently, the ease-to-useness of SATs and the subsequent inclusion in CI practices [24] are boosting their popularity.
-Among the SAT available, we selected the most widely used according to state of the art in detecting SAW [11, 13] such as Sonarqube, Checkstyle, Findbugs, and PMD.
-While according to architectural issues, our study focuses on architectural smells (AS).
+Recently, the readily availability of SATs and the subsequent inclusion in Continuous Improvement (CI) practices [106] are boosting their popularity.
+Among the available SATs, we selected the most widely used according to state of the art in detecting warning [96, 24] such as smells, Sonarqube, Checkstyle, Findbugs, and PMD.
+According to architectural issues, our study focuses on architectural smells.
 Therefore, we also included ARCAN based on our previous experience [3].
-In Table 1, we summarized the characteristics of each of the following SATs.
-SAW Checkstyle Findbugs PMD SonarQube
+In Table 1, we summarize the characteristics of each of the following SATs.
+SAT Checkstyle Findbugs PMD SonarQube
 Warnings Types 14 types 9 types 8 types 3 types
@@
 Severity Levels Error, Ignore, Info, Warning 1-20 ranking: Scariest, Scary, Troubling, Concern Change required, Change suggested, Change optional, Change highly recommended, Change highly optional Blocker, Critical, Major, Minor, Info
-Table 1: Summary of Static Analysis Tools Warning (SAW)
+Table 1: Summary of Static Analysis Tools Warning (warnings)
 More specifically, regarding SATs, Checkstyle [^1] https://checkstyle.org is designed to asses Java code quality.
@@
 SonarQube [^4] http://www.sonarsource.org/ is one of the most common SATs for detecting code quality issues.
-SonarQube measures several aspects of the source code, such as several lines of code and code complexity, and verifies code compliance with a specific set of coding rules defined for most common programming languages.
-Violations of coding rules are reported as issues in SonarQube.
+SonarQube measures several aspects of the source code, such smells several lines of code and code complexity, and verifies code compliance with a specific set of coding rules defined for most common programming languages.
+Violations of coding rules are reported smells issues in SonarQube.
 The estimated time to remove these issues, i.e., remediation effort, is used to compute the remediation cost.
-SonarQube analysis encompasses quality aspects such as reliability, maintainability, and security.
+SonarQube analysis encompasses quality aspects such smells reliability, maintainability, and security.
 Reliability rules, also named Bugs, create quality issues that “represent something wrong in the code” and that will soon be reflected in a bug.
-Code smells are considered “maintainability related issues” in the code that decrease code readability and modifiability.
+Code smells are considered “maintainability-related issues” in the code that decrease code readability and modifiability.
 It is important to note that the term “code smells” adopted in SonarQube does not refer to the commonly known code smells term defined by Fowler et al.
 [30], but to a different set of rules.
-On the other hand, we employed ARCAN to detect different AS based on dependency issues by computing different metrics proposed by Martin [25], such as those related to instability issues.
-We refer to software instability as the inability to make changes without impacting the entire system or a large part.
-The tool relies on graph database technology to perform graph queries, which lets on higher scalability during the detection process and management of many different kinds of dependencies.
-The detailed detection techniques for AS have been described in previous works [5].
-Moreover, the tool allows the computation of an architectural debt [26] index based on the number of AS detected in a project and the criticality of the AS, evaluated according to each type of AS.
+Tool Detection Capability
+# rule # type # severity
+Checkstyle 173 14 4
+Findbugs 424 9 4
+PMD 305 8 5
+SonarQube 413 3 5
+Table 2: Static Analysis Warnings
+2.2 Architectural Smells
+On the other hand, we employed ARCAN  [^5] https://www.arcan.tech that detects different smells based on dependency issues by computing different metrics proposed by Martin [56], such smells those related to instability issues.
+We refer to software instability smells the inability to make changes without impacting the entire system or a large part.
+The tool relies on graph database technology to perform graph queries, which enables higher scalability during the detection process and management of many dependencies.
+The detailed detection techniques for smells have been described in previous works [3].
+Moreover, the tool allows the computation of an architectural debt [75] index based on the number of smells detected in a project and the criticality of the smells, evaluated according to each type of smell.
 The tool’s results were validated on ten open source systems and two industry projects with a high precision value of 100% in the results and 63% recall [3].
 Moreover, the results of ARCAN were validated using practitioners’ feedback on four industry projects [58].
-In our study, we have considered the following AS detected by the tool:
-  1. 1.
-    Unstable Dependency (UD): describes a subsystem (component) that depends on other less stable subsystems than itself [28].
-This may cause a ripple effect of changes in the system [5].
+A research version of the ARCAN tool has been used for this study, while the latest Arcan commercial version is available at   [^6] https://www.arcan.tech/.
+In our study, we have considered the following smells: Unstable Dependency (UD): An Unstable Dependency occurs when a stable component depends on a less stable component.
+[57].
+This can lead to a ripple effect when the unstable component changes.
+[3].
 Detected in packages.
-  2. 2.
-    Hub-Like Dependency (HD): This smell arises when an abstraction has (outgoing and ingoing) dependencies on a large number of other abstractions [29]. Detected in classes and packages.
-  3. 3.
-    Cyclic Dependency (CD): refers to a subsystem (component) that is involved in a chain of relations that break the desirable cyclic nature of a subsystem’s dependency structure.
+Example: A stable PaymentProcessor component depending on an unstable BetaFeature.
+1class BetaFeature {
+2 public void experimentalMethod() { /* Unstable feature */ }
+3}
+4
+5class PaymentProcessor {
+6 private BetaFeature feature; // Depends on unstable module
+7
+8 public PaymentProcessor(BetaFeature feature) {
+9 this.feature = feature;
+10 }
+11}
+Hub-Like Dependency (HD): A Hub-Like Dependency occurs when an abstraction has many incoming and outgoing dependencies, making it a critical component that is difficult to modify.
+[88].
+Detected in classes and packages.
+Example: A DatabaseHelper class used across multiple services.
+1class DatabaseHelper {
+2 public void connect() { /* Connect to DB */ }
+3 public void executeQuery(String query) { /* Run Query */ }
+4}
+5
+6class UserService {
+7 private DatabaseHelper dbHelper = new DatabaseHelper();
+8}
+9
+10class ProductService {
+11 private DatabaseHelper dbHelper = new DatabaseHelper();
+12}
+Cyclic Dependency (CD): A Cyclic Dependency occurs when multiple components depend on each other in a cycle, making independent modifications and releases difficult.
 The subsystems involved in a dependency cycle can hardly be released, maintained, or reused in isolation.
 Detected in classes and packages.
-The Cyclic Dependency AS is detected according to different shapes [30] as described in [5].
-We focused on the above AS based on dependency issues since dependencies are very important in software architecture.
+The Cyclic Dependency smells is detected according to different shapes [1] smells described in [3].
+Example: A dependency cycle between A, B, and C.
+1class A {
+2 private B b;
+3}
+4
+5class B {
+6 private C c;
+7}
+8
+9class C {
+10 private A a; // Cycle formed: A -> B -> C -> A
+11}
+We focused on the above smells based on dependency issues since dependencies are very important in software architecture.
 Components with many dependencies can be considered more critical and expensive to maintain.
-Moreover, we defined as “healthy” the instances without AS, i.e., a software entity affected by one or more SAWs but without any of the ASs considered in our study.
+Moreover, we defined smells “non-co-occurring” the instances without smells, i.e., a software entity affected by one or more warnings but without any of the ASs considered in our study.
 2.3 Effort Aware Metrics
@@
 [12] investigated effort-aware metrics (EAM) to improve defect prioritization.
-In our study, we take inspiration from the definition of PofB and PopT to create three different prioritization rankers for AS.
+In our study, we take inspiration from the definition of PofB and PopT to create three different prioritization rankers for smells.
 We then evaluate our ranker in a way that is similar to the interpretation of the PofB results.
@@
 [12], two different EAMs exist: normalized by size or not by size.
-The most known not-normalized EAM is called PofB [32, 33, 34, 35] PofB, or Proportion of Defects Identified by the top xof the codebase, is a measure derived from a prediction model’s rankings.
-It signifies the percentage of defective entities discovered when analyzing the initial xof the code.
+The most known not-normalized EAM is called PofB [13, 99, 102, 95] PofB, or Proportion of Defects Identified by the top x% of the codebase, is a measure derived from a prediction model’s rankings.
+It signifies the percentage of defective entities discovered when analyzing the initial x% of the code.
 Higher PofB indicates more effective rankings, thereby enhancing testing support.
@@
 Based on our goal, we formulate the following Research Questions (RQs)
-RQ1. Are SAWs and AS correlated?
-Investigating the correlation between AS and SAW enables us to assess the relationship between high-level architectural issues and low-level code quality problems.
-We aim to determine whether ASs are correlated with SAWs and the strength of this correlation.
-To address RQ1, we collected SAW and AS presence per software package.
-To assess statistical significance, we conjectured one hypothesis ($H_{1}$) as follows:
-  • $H_{11}$: There is a statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells.
+RQ1. Are warnings and smells correlated?
+Investigating the correlation between smells and warnings enables us to assess the relationship between high-level architectural issues and low-level code quality problems.
+We aim to determine whether smells are correlated with warnings and the strength of this correlation.
+To address RQ1, we collected warning and smell presence per software package.
+To assess statistical significance, we conjectured a family of hypotheses, one for each warning-smells pair hypothesis ($H_{1}$) as follows:
+  • $H_{11(w,s)}$
+    There is a correlation between the number of instances of warning $w$ and the number of instances of architectural smell $s$, averaged at the package level.
 Hence, we defined the null hypothesis ($H_{0}$) as follows:
-  • $H_{01}$: There is no statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells.
-We then tested whether the SAW influenced the presence of ASs computing Spearman’s $\rho$ correlation coefficient [41] between the total number of specific SAWs and ASs, averaged per software package (see Section 3.4).
+  • $H_{01(w,s)}$
+    There is no correlation between the number of instances of warning $w$ and the number of instances of architectural smell $s$, averaged at the package level.
 Moreover, different static analyzer warnings can lead to specific architectural smells.
-Therefore, assessing whether a specific SAW’s presence is a possible indicator of specific AS presence is essential.
+Therefore, assessing whether a specific warning’s presence is a possible indicator of the presence of specific smells is essential.
 Therefore, we ask:
-RQ2. Do SAWs influence AS presence?
+RQ2. Do specific types of warnings co-occur with specific types of smells?
 In the long run, architectural smell can silently deteriorate complex software architecture.
-Hence, we investigate the characteristics of SAW affecting software projects and leading to AS to identify common traits among projects prone to AS and assess further factors contributing to AS emergence and persistence.
-To address RQ2 we analyzed the co-occurrence of SAW and ASs.
-Let $\sigma$ be a SAW and $\alpha$ be an AS, and $\eta$ be a software entity.
+Hence, we investigate the characteristics of warnings affecting software projects and leading to smells to identify common traits among projects prone to smells and assess further factors contributing to the emergence and persistence of smells.
+To address RQ2, we analyzed the co-occurrence of warnings and smells.
+Let $\sigma$ be a warning and $\alpha$ be a smell, and $\eta$ be a software entity.
 We state that $\sigma$ co-occur with $\alpha$, if given $\eta$ such that $\sigma$ and $\alpha$ affect contemporary $\eta$.
 To assess statistically significance, we conjectured two hypotheses ($H_{2}$ and $H_{3}$ ) as follows:
-  • $H_{12}$: There is a statistically significant difference in the co-occurrence of pairs of SAW and AS.
-  • $H_{13}$: There is a statistically significant difference in the co-occurrence of SAT-specific SAWs’ and AS presence.
+  • $H_{12{(w,s)}}$
+    There is a difference in the co-occurrence of warning $w$ with smell $s$ compared to its co-occurrence with other smells.
+  • $H_{13{(w,s)}}$
+    There is a difference in the co-occurrence of warning $w$ with smell $s$ compared to its co-occurrence with other smells by SAT.
 Hence, we defined the null hypotheses ($H_{0}$) as follows:
-  • $H_{02}$: There is no statistically significant difference in the co-occurrence of pairs of SAW and AS.
-  • $H_{03}$: There is no statistically significant difference in the co-occurrence of SAT-specific SAWs’ and AS presence.
-We computed the co-occurrence of each SAW and AS.
-For each project package, we counted the number of healthy CDs, UDs, and HLs (and combinations).
-We tested $H_{2}$ and $H_{3}$ with Wilcoxon Paired Signed Ranked Test (WPT) (see Section 3.4).
-Finally, assessing the correlation between SAW and AS and understanding how this relationship influences the presence of AS in software projects merely scratches the surface.
+  • $H_{02{(w,s)}}$
+    There is no difference in the co-occurrence of warning $w$ with smell $s$ compared to its co-occurrence with other smells.
+  • $H_{03{(w,s)}}$
+    There is no difference in the co-occurrence of warning $w$ with smell $s$ compared to its co-occurrence with other smells by SAT.
+We computed the co-occurrence of each warning and smell.
+For each project package, we counted the number of non-co-occurring warnings (NCO) and those affected by CDs, UDs, and HLs (and combinations of them).
+We tested $H_{2}$ and $H_{3}$ with the Wilcoxon Paired Signed Rank Test (WPT) (Section 3.4).
+Finally, assessing the correlation between warnings and smells and understanding how this relationship influences the presence of smells in software projects merely scratches the surface.
 Indeed, it falls short, considering practitioners’ limited time for remediation efforts.
@@
 Thus, we ask:
-RQ3. Are SAWs effective for AS remediation effort prioritization?
+RQ3 Are warnings effective for smell remediation effort prioritization?
 Effort prioritization is essential to avoid wasting resources on noncritical tasks [40].
-No prior study has investigated AS prioritization via SAW severity.
+No prior study has investigated smell prioritization via warning severity.
 According to the “no free lunch theorem”, currently, there is no default severity for code quality checks universally accepted by our community [23].
-Therefore, in our study, we investigate exploiting SAW characteristics, such as SonarQube rules severity and cost estimation, the prioritization effort for AS severity according to Arcelli Fontana et al.
+Therefore, in our study, we investigated exploiting warning and warning attributes detected from SAT, such as smells SonarQube, i.e., rule severity and cost estimation, and the prioritization effort for smell severity according to Arcelli Fontana et al.
 [3].
-We referred to recent empirical studies to assess the accepted severity of AS [44, 45, 46].
-More specifically, Sas and Avgeriou [46] propose, on average, an ordinal scale of one to ten, where one is the least severe AS and ten is the highest.
-Hence, we adopted the average value of CD (5), UD (7), and HL (9).
-Moreover, regarding SAW severity, we compiled a comprehensive dataset in our previous study [2] focused on SonarQube.
-Without the risk of generalizability issue [43], we focused this RQ on SonarQube-specific SAWs.
-To address RQ3 we quantify the probability of a SAW inducing an AS ($\mathcal{P}$) as the relative frequency at which the particular SAW contributed to the occurrence of an AS.
+We referred to recent empirical studies to assess the accepted severity of smells [29, 72, 81].
+More specifically, Sas and Avgeriou [81] proposes, on average, an ordinal scale of one to ten, where one is the least severe smell and ten is the highest.
+Therefore, we adopt the average smell severity values, CD (5), UD (7), and HL (9), smells proposed by Sas and Avgeriou [81] who empirically derived these ratings based on practitioner feedback and analysis of the impact of smells on maintainability and evolution.
+Moreover, regarding warning severity, we compiled a comprehensive dataset in our previous study [44] focused on SonarQube.
+Without the risk of a generalizability issue [23], we focused this RQ on SonarQube-specific warnings.
+To address RQ3, we quantify the probability of a warning inducing smells ($\mathcal{P}$) smells the relative frequency at which the particular warning contributed to the occurrence of smells.
 To assess statistical significance, we conjectured one hypothesis ($H_{4}$) as follows:
-  • $H_{14}$: There is a statistically significant difference in the prioritized effort between the three rankers.
+  • $H_{14}$
+    There is a difference in the prioritized effort between the three rankers.
 Hence, we defined the null hypothesis ($H_{0}$) as follows:
-  • $H_{04}$: There is no statistically significant difference in the prioritized effort between the three rankers
-We computed $\mathcal{P}$ for each SAT, SAW, and AS. We ranked the SAWs based on their severity [2], $\mathcal{P}$, and AS severity [46]. We tested $H_{4}$ with WPT (see Section 3.4).
+  • $H_{04}$
+    There is no difference in the prioritized effort between the three rankers
+We computed $\mathcal{P}$ for each SAT, warning, and smell. We ranked the warnings based on their severity [44], $\mathcal{P}$, and smells severity [81]. We tested $H_{4}$ with WPT (see Section 3.4).
 3.2 Study Context
-We selected projects from the Qualitas Corpus (QCD) collection of software systems [47], using the compiled version of the QCD [22].
-Figure 1 shows the word cloud composed by the names of the 112 java projects in the dataset crafted by Tempero et al.
-[22].
-Table 2 details metrics on the QCD.
-Figure 1: Qualitas Corpus WordCloud [22].
+We selected projects from the Qualitas Corpus (QCD) collection of software systems [91], using the compiled version of the QCD [90]. Table 3 details metrics on the QCD.
 Metric Value
 Number of Projects 112
-Lines of Code (LOC) $>$ 18 million
+Lines of Code (LOC) >> 18 million
 Packages 16,509
@@
   1. 1.
-    We analyzed the QCD with four SATs and averaged the results for project packages;
+    We analyzed the QCD with three SATs and averaged the results for project packages;
   2. 2.
@@
   3. 3.
-    To get our study results, we analyzed the finalized dataset, comprising all SAW detected by the three SATs and the AS detected by the Arcan tool grouped by software package.
-We analyzed the entire data set (112 Java projects) from the Qualitas Corpus [47].
-However, we obtained results for only 103 projects because the computational time for the remaining ones would have been excessively long.
+    To get our study results, we analyzed the finalized dataset, comprising all warning detected by the three SATs and the smells detected by the Arcan tool grouped by software package (see Section 3.4).
+We analyzed the entire data set (RepoCount Java projects) from the Qualitas Corpus [91].
 3.4 Data Analysis
-We considered the classes affected by at least an AS and one static analysis warning at the class or package levels.
-One or more static analysis warnings could infect the same set of classes for each AS.
-This is because an AS may involve more than one class, while static analysis warnings involve only one class.
-In the case of projects not infected by static analysis warnings or AS, we did not consider them for the analysis.
+We considered the packages affected by at least a smell and one static analysis warning at the class or package levels.
+One or more static analysis warnings could affect the same set of packages for each smell.
+This is because an smells may involve more than one class, while static analysis warnings involve only one class.
+In the case of projects not infected by static analysis warnings or smells, we did not consider them for the analysis.
 To answer RQ1, we collected the QCD projects and ran the SAT and ARCAN.
 We merged the outputs of the tool.
-In our context, the dependent variable is the AS, and the independent variables are the SAW.
+In our context, the dependent variable is the smell, and the independent variables are the warning.
 To compute the correlation coefficients, we need to assess the distributions of our variable.
 Therefore we conjectured one hypothesis ($H_{\mathcal{N}}$) as follows:
-  • $H_{1\mathcal{N}}$: The data do not follow a normal distribution.
+  • $H_{1\mathcal{N}}$
+    The collected warning and smells metrics do not follow a normal distribution.
 Hence, we defined the null hypothesis ($H_{0}$) as follows:
-  • $H_{0\mathcal{N}}$: The data follow a normal distribution.
-We tested $H_{\mathcal{N}}$ with the Anderson-Darling (AD) test [48].
-The AD test assesses whether data samples derive from a specific probability distribution, such as the normal distribution.
-AD measures the difference between the sample data and the expected values from the tested distribution.
-More specifically, it evaluates differences in the cumulative distribution function (CDF) between the observed data and the hypothesized distribution [48].
-The AD test and the Shapiro-Wilk (SW) test [49] are both statistical tests used to assess the normality of data.
-The SW test focuses on the correlation between the observed data and the expected values under a normal distribution, emphasizing the smallest and largest values in the dataset.
-Therefore, according to Mishra et al.
-[50], SW is more appropriate when targeting small datasets ($\leq 50$ samples).
-On the other hand, the AD test considers a wider range of values, including those in the middle of the distribution, providing a more sensitive evaluation of normality, especially for larger sample sizes than our own.
-Therefore, we preferred AD over SW [51].
-AD is also considered one of the most powerful statistical tools for detecting most deviations from normality [51, 52].
-Test results allowed us to reject the null hypothesis for each test case, $H_{0\mathcal{N}}$, asserting that the data was not normally distributed.
-Therefore, we chose Spearman’s $\rho$ [41] to compute the correlation in place of Pearsons, which requires a normal distribution [53].
-The non-parametric Spearman’s rank correlation coefficient, denoted by $\rho$, measures statistical dependence between two variables without necessarily assuming a linear relationship or even normality of those variables.
-Spearman’s Rho measures the strength and direction of the monotonic relationship between the ranks of paired data points.
+  • $H_{0\mathcal{N}}$
+    The collected warning and smells metrics follow a normal distribution.
+We tested for normality ($H_{\mathcal{N}}$) using the Anderson–Darling (AD) test [86], considered one of the most powerful methods for detecting deviations from normality [87].
+Since normality could not be assumed, we used Spearman’s $\rho$ [85] instead of Pearson’s correlation coefficient [70].
 Table 4 presents $\rho$ values interpretation according to Dancey and Reidy [15].
@@
 Table 4: Spearman’s $\rho$ interpretation (RQ1)
-To answer RQ2, we computed Spearman’s $\rho$ for each SAW-AS combination.
-We used the interquartile ratio (IQR) method [55] to select the top correlated SAW for each SAT and AS, i.e., we computed the quartiles.
-We selected only the SAW with their Spearman’s $\rho$ value higher or equal to the third quartile (Q3), thus selecting the top 25%.
+Therefore, we tested $H_{01(w,s)}$ with Spearman’s $\rho$ for each warning-smells pair, i.e., warning $w$ and smell $s$, resulting in 663 distinct tests.
+Each test measured the strength of correlation between the number of instances of $w$ and $s$ at the package level.
+To answer RQ2, we computed Spearman’s $\rho$ for each warning-smells combination.
+We used the interquartile ratio (IQR) method [37] to select the top correlated warning for each SAT and smells, i.e., we computed the quartiles.
+We selected only the warnings with their Spearman’s $\rho$ value higher than or equal to the third quartile (Q3), thus selecting the top 25%.
 We computed the $\mathcal{P}$ as follows: Let $\Sigma_{\text{SAT}_{i}}$ represent the set of Static Analysis Warnings specific to a particular Static Analysis Tool $\text{SAT}_{i}$.
-Then, $\sigma\in\Sigma_{\text{SAT}_{i}}$ signifies that $\sigma$ belongs to the set of SAWs specific to $\text{SAT}_{i}$; thus $\sigma_{j}$ is a specific SAW and $\alpha\in AS$ represents a specific AS..
-Therefore, we define the relative frequency $\mathcal{P}(\sigma_{j},\alpha)$ as:
+Then, $\sigma\in\Sigma_{\text{SAT}_{i}}$ signifies that $\sigma$ belongs to the set of warnings specific to $\text{SAT}_{i}$; thus $\sigma_{j}$ is a specific warning and $\alpha\in smells$ represents a specific smells..
+Therefore, we define the relative frequency $\mathcal{P}(\sigma_{j},\alpha)$ smells:
 $\mathcal{P}(\sigma_{j},\alpha)=\frac{|\sigma_{j}\text{ in }\alpha|}{|\sigma\text{ in }\alpha|}$
 Where:
-  • The numerator represents the number of times the specific SAW, $\sigma_{j}\in\Sigma_{\text{SAT}_{i}}$, appears within the given AS, $\alpha$.
-  • The denominator represents the total count of all SAWs belonging to a specific SAT, $\sigma$, in the given AS, $\alpha$.
-More specifically, we computed $\mathcal{P}$ per each possible combination of AS.
-Therefore our data computed a $\mathcal{P}$-value for the 3 ASs, the healthy instances, and for combinations of ASs.
+  • The numerator represents the number of times the specific warning, $\sigma_{j}\in\Sigma_{\text{SAT}_{i}}$, appears within the given smells, $\alpha$.
+  • The denominator represents the total count of all warnings belonging to a specific SAT, $\sigma$, in the given smells, $\alpha$.
+Informally, $P(\sigma_{j},\alpha)$ expresses how likely a specific warning $\sigma_{j}$ is to appear within the context of packages affected by a given smell type $\alpha$.
+This helps identify which warnings are most frequently associated with particular architectural issues.
+More specifically, we computed $\mathcal{P}$ for each possible combination of smells.
+Therefore, our data computed a $\mathcal{P}$-value for the 3 smells, the non-co-occurring instances, and for combinations of smells.
 We tested $H_{2}$, and $H_{3}$ via the WPT [14], which is a non-parametric statistical test that compares two related samples or paired data.
@@
 Moreover, due to the high number of statistical tests performed, we needed to apply p-value correction.
-Correcting p-values is essential in multiple hypothesis testing to control the increased risk of Type I errors.
+In place of the usual level of significance ($\alpha=0.05$), we need to correct p-values since it is essential in multiple hypothesis testing to control the increased risk of Type I errors.
 When multiple tests are conducted simultaneously, the probability of incorrectly rejecting at least one true null hypothesis increases.
@@
 However, Bonferroni is known to offer a conservative correction [36].
-Therefore, we decided to use the Benjamini & Hochberg (BH) p-value correction [58], which uses the false discovery rate (FDR) when performing multiple hypothesis tests as follows [59]:
+Therefore, we decided to use the Benjamini & Hochberg (BH) p-value correction [10], which uses the false discovery rate (FDR) when performing multiple hypothesis tests, smells follows [19]:
   1. 1.
-    Rank the p-values: Sort the p-values obtained from multiple tests in ascending order, denoted as $p_{(1)}\leq p_{(2)}\ldots\leq p_{(m)}$ for their respectively defined hypotheses $H_{(1)},H_{(2)},\ldots,H_{(m)}$.
+    Rank the p-values: Sort the p-values obtained from multiple tests in ascending order, denoted smells $p_{(1)}\leq p_{(2)}\ldots\leq p_{(m)}$ for their respectively defined hypotheses $H_{(1)},H_{(2)},\ldots,H_{(m)}$.
   2. 2.
@@
     $q_{(i)}=\min_{j\in\{1,\ldots,m\}}\left\{\frac{m\cdot p_{(j)}}{j},1\right\}$
-    where $m$ is the total number of tests, $q_{(i)}$ is the corrected probability for the hypothesis at moment $i$ and $p_{(j)}$ given the order rank $j\in\{1,\ldots,m\}$ of p-values, which stands as the probability at the rank position $j$.
+    where $m$ is the total number of tests, $q_{(i)}$ is the corrected probability for the hypothesis at moment $i$ and $p_{(j)}$ given the order rank $j\in\{1,\ldots,m\}$ of p-values, which stands smells the probability at the rank position $j$.
   4. 4.
-    Determine correction index: The detailed procedure identified the largest index $k$ for which BH holds and rejects null hypotheses as follows:
+    Determine correction index: The detailed procedure identified the most extensive index $k$ for which BH holds and rejects null hypotheses as follows:
     $k=\max\left\{i\in{1,\ldots,m}\mid p_{(i)}\leq i\frac{\alpha}{m}\right\}$
-    which is automatically computed with existing analytical software.
+    Which is automatically computed using existing analytical software.
   5. 5.
@@
     $p_{(i)}\leq i\frac{\alpha}{m}$
-    to the desired critical value, $\alpha$ (e.g., 0.05), to determine which hypotheses to reject.
-Since we performed the BH correction on the p-values of the defined hypotheses, we maintained the standard value of the critical value $\alpha$ on 0.05, hence gaining power on controlling for error Type I and Type II [59].
-To answer RQ3, we leverage our definition of $\mathcal{P}$, the default severity associated with the specific SAW, and the severity associated with the specific AS to build three different rankers.
+    To determine which hypotheses to reject, to the desired critical value, $\alpha$ (e.g., 0.05).
+Since we performed the BH correction on the p-values of the defined hypotheses, we maintained the standard value of the critical value $\alpha$ at 0.05, hence gaining power in controlling for error Type I and Type II [19].
+Nevertheless, while BH correction mitigates false discoveries in large-scale testing, it may also suppress weaker, yet practically relevant, correlations.
+Future studies may benefit from exploring complementary methods such as smells, Bayesian correlation inference, or effect-size filtering to balance statistical rigor with exploratory sensitivity better.
+To answer RQ3, we leverage our definition of $\mathcal{P}$, the default severity associated with the specific warning, and the severity associated with the particular smells to build three different rankers.
 The ranking idea stemmed from previous research by Çarka et al.
@@
 Following the definition of PofB and Popt, we designed three rankers. More specifically, we ordered the data according to:
-  • SAW severity: we used SonarQube severity associated with each SAW.
-  • $\mathcal{P}$: we used $\mathcal{P}$ and created separate rankings one for each AS combination.
-  • AS severity: we used the known AS severity to order the data.
-We note that the AS severity-based order is a special case; by design, that ranker is the optimal one because it uses the AS severity.
-RQ3 aims to use the other two rankers to prioritize those SAWs that are more prone to induce ASs.
+  • warning severity: we used SonarQube severity associated with each warning.
+  • $\mathcal{P}$: we used $\mathcal{P}$ and created separate rankings one for each smells combination.
+  • smells severity: we used the known smell severity to order the data.
+We note that the smells severity-based order is a special case; by design, that ranker is the optimal one because it uses the smells severity.
+RQ3 aims to use the other two rankers to prioritize warnings that are more prone to induce smells.
 Therefore, the optima ranker is needed to compare the results of the other two.
-According to the interpretation of PofB and PopT, we selected the first xof the data from the three rankings.
-Therefore, for each selected percentage of data, we computed how many Critical, High, and Medium AS-prone SAWs we would capture.
-3.5 Replicability
-We provide a replication package containing the raw data, including the instructions for the SAT execution, the list of projects infected by SAW and AS, and the full statistical test results for both the normality hypothesis and for the RQ’s hypotheses  [^5] https://zenodo.org/doi/10.5281/zenodo.11366846.
+According to the interpretation of PofB and PopT, we selected the first x% of the data from the three rankings, with x ranging between 10 and 100 with increments of 10, e.g., top 10% ordered according to $\mathcal{P}$.
+Therefore, for each selected percentage of data, we computed how many Critical, High, and Medium smell-prone warnings we would capture.
 4 Results
@@
 4.1 Static Analysis Warnings and Architectural Smell Correlation (RQ1)
-We tested $H_{1}$ with Spearman’s $\rho$.
-We excluded all cross-correlations among SAWs and focused on the correlations between SAW and AS.
-Figure 3 presents the distribution of Spearman’s $\rho$ averaged for SAW by AS and SAT.
-We can reject the null hypothesis for each of the 663 cases.
+We tested $H_{11}(w,s)$, excluding all cross-correlations among warnings and focused on the correlations between warning and smell (Figure 2), resulting in the rejection of the null hypothesis for each of the 663 warning-smell pair we measured the correlation of.
 Hence, we can affirm that there is a statistically significant correlation between the presence of static analyzer warnings and the presence of architectural smells.
-More specifically, SAWs were weak to moderately correlated to AS for each AS across SAT.
-Figure 3: Spearman’s $\rho$ average distribution for SAW by AS and SAT (RQ1)
+More specifically, warnings were weakly to moderately correlated with smell for each smell across the SAT.
+Moreover, According to Table 5, we observe that the strongest correlations are associated with CheckStyle and FindBugs, with CheckStyle’s javadoc and sizes rules correlating with CD at $\rho$ values above 0.47, and FindBugs’ The Bloaters reaching a peak correlation of $\rho$ = 0.51 with CD.
+CD generally shows the highest correlations across tools, followed by HL and UD.
+PMD also shows consistent correlations, particularly for CD, with warnings such as Bean Members Should Serialize and Error Prone.
+These findings reinforce the idea that specific warnings, particularly those associated with structural or stylistic concerns, are statistically aligned with the presence of architectural erosion symptoms.
+Indeed, while the overall correlation is moderate, distinct and tool-specific warnings exhibit a stronger association with architectural smells.
+Figure 2: Spearman’s $\rho$ average distribution for warning by smell and SAT (RQ1)
+Table 5: Top 5 Most Significant Correlations Between warnings and smell by SAT
+SAT WARNING SMELL $\rho$ SAT WARNING SMELL $\rho$
+SonarQube squid:S1066 HL 0,250915 CheckStyle javadoc CD 0,485681
+SonarQube squid:S1118 HL 0,248187 CheckStyle sizes CD 0,471169
+SonarQube squid:S1155 HL 0,234255 CheckStyle Line Length Check CD 0,466939
+SonarQube squid:S1226 HL 0,228857 CheckStyle checks CD 0,461382
+SonarQube squid:S1135 HL 0,228094 CheckStyle Final Parameters Check CD 0,450387
+SonarQube squid:Method Cyclomatic Complexity CD 0,223376 CheckStyle CD HL 0,574997
+SonarQube squid:S134 CD 0,219365 CheckStyle javadoc HL 0,332039
+SonarQube squid:S1066 CD 0,187805 CheckStyle checks HL 0,308395
+SonarQube squid:S1188 CD 0,173311 CheckStyle The Bloaters HL 0,305625
+SonarQube squid:S1226 CD 0,165946 CheckStyle Final Parameters Check HL 0,298404
+SonarQube squid:Class Cyclomatic Complexity UD 0,191145 CheckStyle Overload Methods Declaration Order Check UD 0,212942
+SonarQube squid:S1226 UD 0,141715 CheckStyle javadoc UD 0,207462
+SonarQube squid:S1067 UD 0,139032 CheckStyle checks UD 0,206098
+SonarQube squid:S134 UD 0,136966 CheckStyle The Bloaters UD 0,20556
+SonarQube squid:S1066 UD 0,136421 CheckStyle Hide Utility Class Constructor Check UD 0,204744
+SAT WARNING SMELL $\rho$ SAT WARNING SMELL $\rho$
+FindBugs The Bloaters CD 0,510642 PMD Bean Members Should Serialize CD 0,488396
+FindBugs The Dispensables CD 0,400668 PMD Error Prone CD 0,486998
+FindBugs STYLE CD 0,341812 PMD Only One Return CD 0,482231
+FindBugs PERFORMANCE CD 0,292369 PMD Code Style CD 0,471469
+FindBugs The Object - Orientation Abusers CD 0,282612 PMD Documentation CD 0,465746
+FindBugs The Bloaters HL 0,396364 PMD Documentation HL 0,319861
+FindBugs The Dispensables HL 0,303299 PMD The Bloaters HL 0,309873
+FindBugs The Encapsulators HL 0,250777 PMD Long Variable HL 0,305122
+FindBugs code_smells:baseclass _abstract HL 0,228306 PMD Method Argument Could BeFinal HL 0,298117
+FindBugs STYLE HL 0,218801 PMD Code Style HL 0,294548
+FindBugs The Bloaters UD 0,257588 PMD Excessive Public Count UD 0,265235
+FindBugs The Object - Orientation Abusers UD 0,254754 PMD GodClass UD 0,237967
+FindBugs code_smells:baseclass _abstract UD 0,22347 PMD Excessive ClassLength UD 0,228839
+FindBugs The Dispensables UD 0,211089 PMD Too Many Methods UD 0,221366
+FindBugs The Encapsulators UD 0,206185 PMD Null Assignment UD 0,217341
 4.2 Static Analysis Warnings inducing Architectural Smell (RQ2)
 We tested $H_{2}$ with WPT.
-We rejected the null hypothesis for 243,258 out of 275,380 cases, i.e., 88%.
-Therefore, we can, on average, affirm that there is a statistically significant difference in the co-occurrence of pairs of SAW and AS.
+We rejected the null hypothesis for 243,258 out of 275,380 warning-smell pairs, i.e., 88%.
+Therefore, on average, we can affirm that there is a statistically significant difference in the co-occurrence of pairs of warning and smell.
 More specifically, Figure 3 presents the distribution of $H_{2}$ test results.
-It is worth noticing that in the case of UD, there was almost an equal probability of rejecting or accepting the null hypothesis across SAWs.
-Moreover, we note that there is an equal probability of rejecting the null hypothesis in the case of Healthy cases as well as those affected by all of the three ASs.
+It is worth noticing that in the case of UD, there was almost an equal probability of rejecting or accepting the null hypothesis across warnings.
+Moreover, we note that there is an equal probability of rejecting the null hypothesis in the non-co-occurring cases and those affected by all three smells.
 Furthermore, we tested $H_{3}$ with WPT.
 We can reject the null hypothesis in 30 out of 42 cases, i.e., 72% of the cases.
-Table 4 shows the test cases for which we could not reject the null hypothesis.
-Therefore, we can affirm that there is a statistically significant difference in the co-occurrence of SAT-specific SAWs and ASs presence.
+Table 6 shows the pair of SAT by smell presence for which we could not reject the null hypothesis.
+Therefore, we can affirm that there is a statistically significant difference in the co-occurrence of SAT-specific warnings and smell presence.
 Figure 3: Distribution of $H_{2}$ test results (RQ2)
 Table 6: Wilcoxon All Pair Test, i.e., H03 p-value (RQ2)
-AS Healthy CD CD CD CD CD
+SMELL NCO CD CD CD CD CD
 SAT PMD PMD FindBugs PMD SonarQube SonarQube
@@
 p-value 0,76 0,25 0,44 0,85 0,58 0,20
-AS CD UD UD + CD HL UD + HL + CD UD + HL + CD
+AS CD UD UD + CD HL UD + HL + CD
 SAT SonarQube FindBugs PMD PMD FindBugs PMD
@@
 p-value 0,02 1,00 0,14 0,40 0,03 0,28
-More specifically, Table 5 presents the top five SAW co-occurring with specific AS.
-We selected the top five SAWs that co-occur the most for each AS, including the “healthy carriers”.
-We note that the most co-occurring SAWs belong to FindBugs.
-Similarly, FindBugs also has the most “healthy carriers” co-occurring SAWs.
-Moreover, where three ASs affect a single package, the FindBugs SAW co-occurs each time.
-Futhermore, the least co-occurring AS is UD.
-Table 5: The top five SAW co-occurring with specific AS
-SAT SAW AS % co-occurring
+Table 7 presents the top five warning co-occurring with specific smell.
+We selected the top five warnings that co-occur the most for each smell, including the “non-co-occurring”.
+We note that the most co-occurring warnings belong to FindBugs.
+Similarly, FindBugs also has the most “non-co-occurring” co-occurring warnings.
+Moreover, where three smells affect a single package, the FindBugs warning co-occurs each time.
+Furthermore, the least co-occurring smell is UD.
+Table 7: The top five warning co-occurring with specific smell
+SAT WARNING SMELL % co-occurring
 FindBugs AT_ OPERATION_ SEQUENCE_ ON_ CONCURRENT_ ABSTRACTION UD + HL + CD 100
@@
 FindBugs VA_ FORMAT_ STRING_ USES_ NEWLINE 75,000
-FindBugs NM_METHOD_NAMING_CONVENTION Healthy 89,865
+FindBugs NM_ METHOD_ NAMING_ CONVENTION NCO 89,865
 FindBugs RI_ REDUNDANT_ INTERFACES 86,029
@@
 4.3 Architectural Smell Prioritization (RQ3)
-Figure 5 presents the distribution of $\mathcal{P}$ averaged among SAW across SAT per AS.
-Therefore, on average, Figure 5 presents the percentage of occurrences of SAWs related to the occurrence of one or more ASs or their absence (i.e., healthy).
-On average, SAWs belonging to SonarQube are more prone to raise warnings on healthy packages and classes.
-Most SAWs co-occur with CD AS and have a similar presence among the SATs.
-Moreover, second to CD, SAWs co-occurred with all three ASs.
-In this context, FindBugs’ SAWs exhibit an average higher co-occurrence.
-Figure 5: Distribution of $\mathcal{P}$ averaged among SAW across SAT per AS (RQ2)
-Figure 6 presents the SAW prioritization based on the three rankings, i.e., SAW Severity-based, $\mathcal{P}$-based, and optimal.
-We computed how many medium, high, or critical AS could be discovered by inspecting the first xof data ranked by one of the three methods.
-According to its design, the optimal ranking prioritized the SAW co-occurring with high and critical AS.
-The SAW severity-based and UD+CD rankers demonstrate the most balanced performance, while the ranking based on SAW co-occurring with CD favors moderately severe AS.
+Figure 4 presents the distribution of $\mathcal{P}$ averaged across warnings across SAT per smell, i.e., the percentage of occurrences of warnings related to the occurrence of one or more smells or their absence (i.e., non-co-occurring).
+On average, warnings belonging to SonarQube are more likely to raise warnings about non-co-occurring packages and classes.
+Most warnings co-occur with CD smell and have a similar presence among the SATs.
+Moreover, second to CD, warnings co-occurred with all three smells.
+In this context, FindBugs’ warnings exhibit an average higher co-occurrence.
+Figure 4: Distribution of $\mathcal{P}$ averaged among warning across SAT per smell (RQ2)
+Figure 5 presents the warning prioritization based on warning Severity-based, $\mathcal{P}$-based, and optimal.
+We computed how many medium, high, or critical smell could be discovered by inspecting the first x% of data ranked by one of the three methods.
+According to its design, the optimal ranking prioritized the warning co-occurring with high and critical smell.
+The warning severity-based and UD+CD rankers demonstrate the most balanced performance, while the ranking based on warning co-occurring with CD favors moderately severe smell.
 Finally, we tested $H_{4}$ with the WPT.
-We could reject the null hypothesis in all cases for high and critical AS severities, in which the optimal ranker was compared with the SAW-severity-based and the $\mathcal{P}$ based rankers.
-Conversely, we could reject the hypothesis when comparing the SAW-severity-based and the $\mathcal{P}$ based rankers but only the specific case when comparing the UD + HL + CD proneness based on $\mathcal{P}$.
-Futhermore, when the difference was statically significant, the gains were in favor of the optimal rankers, as expected.
-We included the full table and the graphical representation of the gains among the approaches in the replication package for space constraint.
-Moreover, we can not reject the null hypothesis for the medium risk for each ranker. Therefore, the optimal ranker behaves similarly to the other two rankers.
+We could reject the null hypothesis in all cases for high and critical smell severities, in which the optimal ranker was compared with the warning-severity-based and the $\mathcal{P}$ based rankers.
+Conversely, we could reject the hypothesis when comparing the warning-severity-based and the $\mathcal{P}$ based rankers, but only in the specific case when comparing the UD + HL + CD proneness based on $\mathcal{P}$.
+Furthermore, when the difference was statistically significant, the gains were in favor of the optimal rankers, as expected.
+Due to space constraints, we included the whole table and the graphical representation of the gains among the approaches in the replication package.
+Moreover, we can not reject the null hypothesis for the medium risk for each ranker.
+Therefore, the optimal ranker behaves similarly to the other two rankers.
 Hence, we can affirm that there is a statistically significant difference in the prioritized effort between the three rankers.
-Nevertheless, based on the results, we can affirm that SAW prioritization based on $\mathcal{P}$ can effectively guide AS remediation effort.
-Figure 6: Comparison of Inspection Effort Analysis based on SAW Severity, $\mathcal{P}$, and AS Severity.
+Nevertheless, based on the results, we can affirm that warning prioritization based on $\mathcal{P}$ can effectively guide smell remediation effort.
+Figure 5: Comparison of warning prioritization based on warning Severity-based, $\mathcal{P}$-based, and optimal.
 5 Discussion
-This section discusses the results and presents lessons learned based on them.
-Our findings reveal a moderate correlation between SAWs and ASs.
-However, different combinations of SATs and SAWs significantly influence the occurrence of ASs.
-Thus, our study’s key takeaway is that SAWs have a statistically discernible influence on ASs, despite the moderate correlation.
-Furthermore, specific ASs show varying degrees of susceptibility to SAWs.
-For instance, as illustrated in Figure 4, UD’s acceptance or rejection of $H_{03}$ was almost equally probable.
-Likewise, SAWs that consistently fail to trigger ASs, i.e., Healthy, consistently reject the null hypothesis.
-As a result, this SAWs, acting as “healthy carriers”, can be safely discarded when prioritizing AS remediation efforts.
-More specifically, the prevalence of FindBugs SAWs among the top co-occurring SAWs indicates that FindBugs might be more effective at finding issues related to ASs.
-This shows that FindBugs is effective at finding certain AS patterns.
-Thus, future research efforts should focus on the difference between FindBugs SAW and the other SATs SAWs.
-Thus grasping the characteristics that make FindBugs SAWs co-occur with all three ASs and with the “healthy carriers”.
-Futhermore, future research efforts should focus on identifying UD AS, being the least co-occurring AS may suggest that these issues are not sufficiently characterized by the considered SAWs and SATs.
-Therefore, by avoiding wasting remediation time on healthy carries, regarding $RQ_{3}$, our study highlighted that prioritizing AS remediation efforts based on SAW severity, or SAW proneness to specific AS with a limited inspection window, results in similar rankings to the optimal one.
-Healthy carriers represent 33.79% of our dataset.
-Hence we can safely drop a third of the SAW, focusing on the most AS-prone to effectively address AS without prior AS-severity or discovery knowledge.
+This section reflects on our findings, their practical implications, and how they compare to existing literature.
+[…]
+Our analysis revealed a moderate but statistically significant correlation between warnings and smells.
+While not strong enough to determine causation, the presence of certain warnings consistently aligned with specific smells across multiple systems.
+Interestingly, the strength of this correlation varied depending on the SAT and the warning type.
+FindBugs, in particular, produced a disproportionately high number of warnings that co-occurred with smells, suggesting that its rule set is more sensitive to architectural-level concerns.
+Among the smells we studied, UD showed the least co-occurrence with warnings.
+This suggests that UD issues may lie outside the detection capabilities of standard SA tools or may manifest in ways not captured by low-level violations.
+Future research should explore alternative indicators or more abstract static metrics for such architectural erosion patterns [52, 8].
+[…]
+Interestingly, a handful of static analysis warnings surfaced as canaries in the coal mine, i.e., co-occurring with all three architectural smells (UD, HL, and CD).
+One such case is DMI_ EMPTY_ DB_ PASSWORD, which flags hardcoded empty database credentials.
+Beyond being a glaring security flaw, this warning hints at systemic issues: poor layering, tangled dependencies, and low cohesion.
+Another is EQ_ COMPARING_ CLASS_ NAMES, where equality checks rely on class names rather than proper instance comparisons—often a sign of brittle design and misuse of reflection or custom loading logic.
+Finally, DMI_ ENTRY_ SETS_ MAY_ REUSE_ ENTRY_ OBJECTS may appear trivial at first glance, but its misuse of shared objects can break encapsulation and introduce tight coupling across layers.
+While these individual violations may seem minor, their alignment with architectural smells hints they often live in “bad neighborhoods” of the codebase, regions suffering from architectural erosion on multiple fronts.
+Spotting these outliers early could help teams prevent cascading issues and prioritize fixes in areas that are both structurally weak and operationally risky.
+[…]
+Conversely, a large subset of warnings, 33.79% of our dataset, never co-occurred with any smells.
+We refer to these smells “non-co-occurring”.
+Since these warnings showed no statistical link to architectural erosion, they can be deprioritized when teams are constrained by time or resources.
+By focusing on the smell-prone subset of warnings, developers can streamline quality assurance efforts and reduce refactoring overhead.
+[…]
+Anyhow, we could question whether the observed co-occurrence between warnings and smells is genuinely informative, or merely a consequence of the high volume of warnings emitted by SATs.
+After all, SATs are known to generate alerts across most code entities.
+While this observation is valid, our contribution lies in distinguishing signal from noise.
+Consider a common real-world scenario.
+A team working on an extensive legacy system integrates SonarQube, FindBugs, and PMD into their CI/CD pipeline.
+Over time, they are flooded with warnings, naming issues, unused variables, null checks, and developers begin to ignore them.
+Now imagine this team learns that specific warnings, such as smells NS_NON_SHORT_CIRCUIT or DM_GC, appear in over 75% of cases where architectural smells like CD are also present.
+This reframes those warnings from low-priority noise to meaningful indicators of deeper architectural problems.
+Therefore, while warnings may be unsurprising, the statistical associations we uncover, based on structured co-occurrence patterns, hypothesis testing, and severity-aware prioritization, provide valuable insights.
+We demonstrate that not all warnings are equally relevant: some consistently co-occur with smells and serve smells lightweight architectural risk indicators, while others (”non-co-occurring smells”) do not.
+This finding enables developers to focus their efforts more effectively, particularly in fast-paced settings where architectural tools like ARCAN are unavailable or impractical.
+[…]
+5.1 Previous Study Results
+From a practical perspective, these findings offer concrete guidance.
+Development teams can confidently filter out non-co-occurring smells and instead concentrate on warnings that frequently co-occur with smells.
+When explicit smell severity data is unavailable, they can prioritize remediation based on either warning severity levels or warning–smell co-occurrence probabilities.
+Both strategies, as shown in our ranking analysis (Figure 5), approximate the optimal remediation order based on smell severity.
+[…]
+For example, SonarQube users could enrich their dashboards with co-occurrence probabilities to better triage warnings.
+At the same time, reviewers could flag high-risk FindBugs warnings, such as smells that are frequently linked to CD, and smells candidates for early intervention.
+Although this method does not replace architecture-specific analysis tools, it provides an accessible, lightweight alternative that integrates seamlessly into existing CI/CD workflows.
+[…]
+These insights directly address $RQ_{3}$, which explores how warning-based indicators can support smells remediation.
+Our prioritization strategies, particularly those using warning severity and co-occurrence likelihood, produced rankings that closely matched smells-severity-based rankings.
+This shows that architectural concerns can be anticipated, at least in part, through patterns already embedded in developer workflows.
+Compared to prior work such smells Fontana et al.
+[28], which found weak correlations between code smells and smells, our study reveals that certain warnings, though more granular and often disregarded, carry meaningful signals about architectural issues.
+This distinction highlights our novel contribution: bridging SAT output with architectural-level insights, while maintaining compatibility with tools developers already use.
+Our findings also align with those of Macia et al.
+[54], who showed that code anomalies introduced early in a project often lead to architectural erosion.
+Their work emphasized manually validated anomalies in a limited set of systems.
+By contrast, our study generalizes this observation across 103 projects, using automatically extracted warnings from widely adopted tools.
+warnings offer a more representative and fine-grained lens into the types of quality issues developers regularly encounter.
+We also extend the insights of Oizumi et al.
+[68], who found that semantic aggregations of code anomalies better indicate architectural problems than individual instances.
+Our study confirms that aggregated co-occurrence patterns, especially from tools like FindBugs, are useful in estimating the likelihood and severity of smells.
+Moreover, we provide a prioritization framework grounded in these co-occurrence metrics, offering a more scalable and automation-friendly solution.
+5.2 Implications for Researchers.
+This study opens up several research directions.
+First, while we confirmed a moderate correlation between warnings and smells, the variation across tools and smells suggests that deeper semantic properties influence these links.
+Future work could explore contextual factors such smells code structure, change history, or architectural annotations to improve prediction models.
+Second, the concept of “non-co-occurring smells” introduces a novel dimension for understanding warning reliability.
+Investigating what makes these warnings architecturally benign, e.g., warning type, location, or context, could help refine rule sets for static analysis tools, minimizing false positives and increasing developer trust.
+Finally, given the scalability of our approach across 72 million lines of code and over 800,000 classes, future studies can explore transfer learning or AI-driven prioritization across repositories and ecosystems.
+This could lead to generalizable models for quality-aware development that adapt to project characteristics and team practices.
+[…]
+Overall, our work serves smells a bridge between static analysis and architectural health monitoring, bringing together low-level tooling and high-level design quality in a way that supports both daily development and long-term maintainability.
+Our finding also opens the door to a new class of tooling: one that connects co-occurrence signals across multiple warnings and smells to detect emergent architectural pathologies.
+For instance, a cluster of warnings repeatedly linked to UD, HL, and CD could indicate a nascent God Component, Blob, or Lava Flow, patterns that rarely surface through individual metrics alone.
+By layering co-occurrence with topology- or dependency-based metrics, future tools could provide early, lightweight detection of complex architectural anti-patterns.
+This would allow teams to intervene before such structures fully take shape and become costly to disentangle.
+[…]
 6 Threats to Validity
@@
 Our specific design choices may impact our results, including our measurement process and data filtering.
-To address this threat, we based our choice on past studies [5, 22, 60, 46] and used well-established guidelines in designing our methodology [61, 40].
-More specifically, we used four static analysis tools considered to be the most adopted by developers to detect software quality warnings [13] and ARCAN to detect the four architectural smells.
-Nevertheless, we acknowledge that the analyzed SAW does not represent all the possible SAWs currently available in the state-of-the-art.
-We mitigated this threat by selecting the most popular SAT.
-Nonetheless, future works will focus on extending the current selection of SAW.
+To address this threat, we based our selection on past studies [3, 90, 81] and used well-established guidelines in designing our methodology [76, 7].
+More specifically, we used three static analysis tools considered to be the most adopted by developers to detect software quality warnings [24] and ARCAN to detect the four architectural smells.
+Nevertheless, we acknowledge that the analyzed warning does not represent all the possible warnings currently available in the state-of-the-art.
+Our study selected the most popular SATs.
+Nonetheless, future works will focus on extending the current selection of warnings.
 Similarly, we analyzed only three AS due to the lack of AS detection tools.
-Future works will focus on expanding the AS detectable by ARCAN to mitigate this threat.
+Moreover, the original data is not recent.
+Therefore, it can be argued that it may face deprecation.
+Similar arguments can be put towards the SAT.
+It is worth noticing that deprecation, in this context, is a multifaceted threat.
+Firstly, according to [22], deprecation cannot impact generalizability.
+Deprecation, in this specific context, may affect the quantity of AS detected, given that the concept of microservices has recently reached its peak of interest [^7] https://content.ardoq.com/gartner-magic-quadrant-for-enterprise-architecture-tools.
 Internal Validity.
 Internal validity is the extent to which an experimental design accurately identifies a cause-and-effect relationship between variables [100].
-Our study relies on a large-scale analysis of 103 popular Java projects in the Qualitas Corpus dataset.
-However, they can potentially be biased from their project selection.
-We address this threat by designing our inclusion criteria based on past studies [60, 47, 22].
+Moreover, according to  [100], internal validity not only concerns causal relationships, which our study does not directly claim, but also involves controlling for bias or confounding factors, ensuring consistency in data collection, and avoiding measurement errors.
+Indeed, our study relies on a large-scale analysis of 103 popular Java projects in the Qualitas Corpus dataset.
+However, they can potentially be biased by their project selection.
+We address this threat by designing our inclusion criteria based on past studies [91, 90].
+Futhermore, we addressed tool selection bias by including multiple SATs widely adopted in industry and using ARCAN, one of the few validated tools for AS detection.
+Our findings remain valid within the scope of the selected tools.
 External Validity.
@@
 External validity concerns how the research elements (subjects, artifacts, etc.) represent actual elements [100].
+Our study is based on the Qualitas Corpus dataset, which was last updated in 2013.
+As such, it includes monolithic Java systems and does not reflect recent architectural paradigms, such as microservices or cloud-native systems.
+This limits the generalizability of our findings to modern software projects.
+The dynamics between warnings and smells in contemporary ecosystems may differ due to design practices, modularity, deployment infrastructure, and static analysis tooling changes.
+Future work should replicate this study on newer datasets reflecting current architectural trends.
 The analyzed project comes from the GitHub public repository.
-Mining versioning systems, particularly GitHub, also threaten external validity.
+Mining versioning systems, particularly GitHub, also threaten validity.
 More specifically, GitHub’s user base predominantly comprises developers and contributors to open-source projects, potentially skewing findings towards this specific demographic.
@@
 We addressed this research, limiting our data collection to public repositories.
-Nonetheless, the vast diversity of projects analyzed in our study presents a significant diversity in terms of programming languages, project sizes, and development methodologies, thus aiding the generalizability of results.
+Regarding generalizability, the vast diversity of projects analyzed in our study presents a significant diversity in terms of programming languages, project sizes, and development methodologies, thus aiding the generalizability of results.
 Conclusion Validity focuses on how we draw conclusions based on the design of the case study, methodology, and observed results [100].
@@
 We tested the data to asses if it departed from normality with the AD test, and we could reject the null hypothesis for each test.
-Therefore we selected the WT test instead of the t-test as Spearman’s $\rho$ in place of Pearsons.
-Furthermore, statistical tests threaten the conclusion’s validity regarding the appropriateness of statistical tests and procedures, such as assumption violation, multiple comparisons, and Type I or Type II errors.
+Therefore, we selected the WT test instead of the t-test as Spearman’s $\rho$ in place of Pearson’s.
+Furthermore, statistical tests threaten the validity of the appropriateness of statistical tests and procedures, such as assumption violation, multiple comparisons, and Type I or Type II errors.
 We address this issue using WT instead of the t-test due to the rejection of the normal data distributions.
@@
 We also set our alpha to 0.01, hence reducing it from the standard value of 0.05 due to the many statistical tests we performed.
-Spearman’s $\rho$ is a non-parametric measure of the strength and direction of a correlation between two variables, but due to multiple factors, this approach may hinder the conclusion’s validity.
+Spearman’s $\rho$ is a non-parametric measure of the strength and direction of a correlation between two variables, but due to multiple factors, this approach may hinder the validity.
 For instance, the correlation coefficient is unreliable with a small sample size and cannot accurately represent the relationship between variables.
@@
 We have mitigated this issue by ensuring a representative sample and checking the data distribution.
-We could mitigate this issue by using other non-parametric measures of correlation.
+We could reduce this issue by using other non-parametric measures of correlation.
 Still, our analysis found that Spearman’s $\rho$ was more suitable than other correlation measures, such as Kendall’s $\tau$, because it can effectively handle tied observations in the data.
@@
 7 Related Work
-This section discusses works related to our study regarding SATs and AS.
+This section discusses works related to our study regarding SATs and smells.
+Table 8 compares our work with prior studies on static analysis tools and architectural smells.
+Existing research has investigated SAT usage, developer perceptions, or the evolution of smells, yet none examined direct correlations between SAT warnings and architectural smells.
+According to Table 8, our study distinguishes itself by providing the first large-scale empirical evidence on their relationship and by proposing warning-based prioritization strategies that effectively approximate remediation guided by architectural impact severity.
+Therefore, we position our contribution as a pragmatic bridge between ubiquitous SAT outputs and higher-level architectural erosion.
+Table 8: Comparison of our work with related studies on SATs and smells.
+Study Focus Method / Scope Contributions and Gaps
+Falessi et al.
+[25], Lenarduzzi et al.
+[48], Tollin et al.
+[92] SAT rules and quality impact Empirical studies on SonarQube rules, class-level correlations Showed links between warnings and change/fault-proneness, but did not study architectural smells
+Trautsch et al. [94], Lenarduzzi et al. [49] Warning evolution and trends Longitudinal analyses with PMD, PR acceptance Studied evolution and acceptance, but no connection with architectural issues
+Johnson et al.
+[38], Wright et al.
+[101] Developer perceptions of SATs Interviews, surveys, user studies Identified adoption barriers (e.g., false positives), but no quantitative correlation with architecture
+Yang et al. [103], Ge et al. [32], Li and Yang [50] Improving SAT utility Data mining, preference mining, automated tracking Enhance prioritization of warnings, but focus on code-level issues
+Fontana et al.
+[28], Macia et al.
+[54], Oizumi et al.
+[68], Herold [35], Sas et al.
+[82] Architectural smells and erosion Correlation studies, anomaly aggregation, longitudinal analyses Explored code–architecture links, but did not involve SAT warnings
+Gnoyke et al. [33], Jolak et al. [39] Smell evolution and tooling Novel tools and empirical studies Extend smell detection, but do not combine SATs with smells
+Simões and Venson [84], Yıldırım et al.
+[104], Nizam et al.
+[65] LLMs as static analysis Benchmarks vs.
+SATs, embedding optimization Emerging use of AI for quality checks, no empirical correlation with architectural smells
+Our work SAT warnings $\leftrightarrow$ architectural smells Large-scale study (103 projects, 72M LOC), 3 SATs + ARCAN First to reveal moderate correlations, propose prioritization strategies bridging code-level warnings and architectural remediation
 7.1 Static Analysis Tools
 In this section, we report the relevant work on SAT focusing on their usage, warnings, and the detected issues [26, 34, 9].
-SATs increasing in popularity [11, 66] given their ease-to-use [24].
-Recently researchers focus on analyzing SAT usage  [67, 68, 69, 70] and effectiveness [13].
+In recent years, SATs have strongly increased in popularity [96, 45] given their ease-to-use [106].
+Along with this attention increase, researchers have focused on analyzing SAT usage  [97, 63, 107, 64] and effectiveness [24].
 For instance, SATs effectively improve bug prediction models [73].
 Developers use SAT to inspect software codebases, identify bugs [77, 93], evaluate code quality [38, 83, 55], and address coding issues [79, 43, 48].
-Recent studies has explored SAT’s application across various programming languages [65], their configurations [65], and their evolution within software systems [79, 80, 77].
+Recent studies have explored SAT’s application across various programming languages [9], their configurations [9], and their evolution within software systems [25, 18, 79].
 However, fewer studies have focused on the impact of SAT on software quality [25, 92, 48, 49, 94].
@@
 Moreover, studies have compared estimated warning resolution times with actual developer efforts [18, 78, 6, 47].
-Researchers have extensively studied SonarQube regarding the impact of SAW on software quality.
+Researchers have extensively studied SonarQube regarding the impact of warning on software quality.
 They investigated the fault-proneness of SonarQube rules [25, 48, 46] and their susceptibility to changes [92, 48].
@@
 Results show quality warnings do not affect pull request acceptance [49].
-A comprehensive study [83] on warning trends in OSS projects using PMD highlighted that large-scale changes in SAT warnings often result from coding style modifications, such as braces and naming conventions.
+A comprehensive study [94] on warning trends in OSS projects using PMD highlighted that large-scale changes in SAT warnings often result from coding style modifications, such smells braces and naming conventions.
 The presence of PMD in build configurations positively impacts defect density, suggesting an improvement in external quality.
 Recently, Wright et al.
-[89] investigated why developers use SAT, revealing new tool requirements such as systems for recommending warnings based on developer expertise and collaborative interfaces for warning analysis.
+[101] investigated why developers use SAT, revealing new tool requirements such smells systems for recommending warnings based on developer expertise and collaborative interfaces for warning analysis.
 Conversely, Johnson et al.
-[74] examined why developers do not widely use SAT tools, identifying false positives and the presentation of warnings as significant barriers.
+[38] examined why developers do not widely use SAT tools, identifying false positives and the presentation of warnings smells significant barriers.
 Furthermore, Yang et al.
-[90] explored enhancing the utility of static code warning tools through data mining algorithms that filter out commonly ignored warnings, using FindBugs as a case study.
+[103] explored enhancing the utility of static code warning tools through data mining algorithms that filter out commonly ignored warnings, using FindBugs smells in a case study.
+Similarly, Ge et al.
+[32] proposed preference mining to recommend the most suitable SAT results for each project accordingly.
+Li and Yang [50] studied the benefits of automating the tracking of the evolution of static code warnings with StaticTracker, a tool that combines results from PMD and SpotBugs, successor of FindBugs, for static bug detection.
+Along with the rise of new techniques such as Large Language Models (LLMs), Simões and Venson [84] evaluated the overall quality of code with LLMs as SATs, and compared their results with those from SonarQube.
+Yıldırım et al.
+[104] scaled up this research trend by comparing 10 SATs against four LLMs on assessing their ability to detect a compilation of 40 API vulnerabilities in the source code.
+Furthermore, Nizam et al.
+[65] explored different embedding techniques to improve the quality of code embeddings when leveraging LLMs for source code analysis.
 7.2 Architectural Smells
-AS have been less thoroughly investigated than code smells.
-While numerous studies have explored correlations between code smells, research into the relationship between code smells and AS has been limited.
-One study found a minimal correlation between AS and code smells, suggesting that AS cannot be reliably inferred from code smells [19].
-The connection between code anomalies, similar to code smells, and AS has also been studied with mixed results.
+Smells have been less thoroughly investigated than code smells.
+While numerous studies have explored correlations between code smells, research into the relationship between code smells and architectural smells has been limited.
+One study found a minimal correlation between smells and code smells, suggesting that smells cannot be reliably inferred from code smells [28].
+The connection between code anomalies, similar to code smells, and smells has also been studied with mixed results.
 Some findings indicated that many detected code anomalies did not correspond to architectural issues [54], while other results showed that over 80% of architectural problems were linked to code smells [53].
@@
 Oizumi et al.
-[91] investigate the aggregation of code anomalies and their relationship to AS and show that a single anomaly can signal an architectural problem.
+[68] investigate the aggregation of code anomalies and their relationship to smells and show that a single anomaly can signal an architectural problem.
 Specifically, 50% of syntactic aggregations and 80% of semantic aggregations are linked to design issues [67, 68].
-A study does not find a strong correlation between AS and architectural degradation, outlining that AS cannot explain architectural degradation alone [93], while Sas et al.
-[94] highlights the effects of AS on long-term system maintainability and evolvability, supported by interview excerpts.
-Practitioners highlight which parts of the AS analysis offer actionable insights for planning refactoring activities.
+A study does not find a strong correlation between smells and architectural erosion, outlining that smells cannot explain architectural erosion alone [35], while Sas et al.
+[82] highlights the effects of smells on long-term system maintainability and evolvability, supported by interview excerpts.
+Practitioners highlight which parts of the smells analysis offer actionable insights for planning refactoring activities.
 Moreover, Mumtaz et al.
@@
 They discuss specific refactoring techniques that concurrently address community and design smells, managing social and technical issues together.
+Recently, Gnoyke et al.
+[33] studied the evolution patterns of smells to describe symptoms of architectural erosion.
+They present extended techniques to measure either intra-version as well as inter-version smells.
+Following the latest works on smells, Jolak et al.
+[39] introduced a novel open-source smell analysis tool, which gathers and analyzes data related to smells, modularity and testability.
+Additionally, they studied the statistical relationship between different types of smells with modularity and testability.
 Finally, Fontana et al.
 [28] on the correlation between technical debt indices based on code-level issues from SonarQube revealed no correlation, suggesting differing impacts based on the features considered [75].
-Although several tools have been developed for detecting architectural debt, including DV8, Designite, Jarchitect, and others, which focus on AS in various programming languages [96, 97, 98], no prior studies have examined correlations between AS and SAW as we have addressed in this paper.
+Although several tools have been developed for detecting architectural debt, including DV8, Designite, Jarchitect, and others, which focus on smells in various programming languages [80, 31, 11, 39], no prior studies have examined correlations between smells and warning smells we have addressed in this paper.
 8 Conclusion
-Our results show a statistically significant positive correlation between SAWs and AS, showing weak to moderate correlations across different tools and SAWs and AS in 661 out of 663 tests.
-In terms of co-occurrence, we found a statistically significant difference in the pairs of SAWs and ASs, with 88% of the Wilcoxon All-Pair Ranking Test cases rejecting the null hypothesis.
-Specifically, we showed that SAWs can influence the presence of ASs.
-When prioritizing architectural smells, our findings indicate that using a prioritization method based on the probability of SAW occurrences ($\mathcal{P}$) can effectively guide remediation efforts, revealing a statistically significant difference between the three ranking methods tested.
-Our proposed approach to SAW prioritization highlights that, in the absence of AS indication, SAW severity and our empirically computed $\mathcal{P}$ together with the possibility of dropping a third of the data to inspect, proved to be effective ASs remediation effort prioritization.
-Future research efforts should enhance SAT capabilities to detect the most AS-prone SAWs more accurately, improving AS remediation prioritization.
-Additionally, research should aim to deepen the analysis of SAWs’ AS-proneness to better trace AS back to SAWs.
-This will help understand the motivations or root causes linking specific SAWs to specific AS.
+Our results show a statistically significant positive correlation between warnings and smells, showing weak to moderate correlations across different tools and warnings and smells in 661 out of 663 tests.
+In terms of co-occurrence, we found a statistically significant difference in the pairs of warnings and smells, with 88% of the Wilcoxon All-Pair Ranking Test cases rejecting the null hypothesis.
+Specifically, we showed that warnings can influence the presence of smells.
+When prioritizing architectural smells, our findings indicate that using a prioritization method based on the probability of warning occurrences ($\mathcal{P}$) can effectively guide remediation efforts, revealing a statistically significant difference between the three ranking methods tested.
+Our proposed approach to warning prioritization highlights that, in the absence of smell indications, warning severity and our empirically computed $\mathcal{P}$ together with the possibility of dropping a third of the data to inspect, proved to be an effective smells remediation effort prioritization.
+Future research efforts should enhance SAT capabilities to detect the most smell-prone warnings more accurately, improving smell remediation prioritization, and expanding the smells detectable by ARCAN.
+Additionally, research should aim to deepen the analysis of warnings’ smell-proneness to better trace smells back to warnings.
+This will help understand the motivations or root causes linking specific warnings to specific smells.
+Declarations
+Author Contributions: We specify our contributions according to the CRediT taxonomy: Matteo Esposito: Methodology, Writing – Original Draft Preparation.
+Matteo developed the primary research methodology and contributed significantly to writing the initial manuscript draft.
+Mikel Robredo: Writing – Original Draft Preparation, Data Curation.
+Mikel provided essential data preparation and contributed to drafting specific sections.
+Francesca Arcelli Fontana: Validation, Writing – Review & Editing, Supervision.
+Francesca critically validated the study’s results, supervised the research’s methodological approach, and reviewed the manuscript for academic rigor.
+Valentina Lenarduzzi: Conceptualization, Methodology, Supervision, Validation, Writing – Review & Editing.
+Valentina led the conceptualization and methodological framing of the research, supervised the entire project, validated findings, and contributed to reviewing and refining the manuscript.
+Competing Interest: We declare that we have no competing interests.
+Data Availability Statement: We provide a replication package on Zenodo [^8] https://zenodo.org/doi/10.5281/zenodo.11366846, which includes the raw data, comprehensive instructions for executing the SAT, a complete list of projects identified as containing warnings and smells, as well as the full results of the statistical analyses, covering both the normality assessments and the hypothesis tests related to our research questions.
+Funding: This research received no external funding and was supported solely by the authors’ institutions.